### v1

# Md Kowsher School of Engineering and Science

Abstract
In Chaos, a minor divergence between two initial conditions exhibits exponential amplification over time, leading to far-away outcomes, known as the butterfly effect. Thus, the distant future is full of uncertainty and hard to forecast. We introduce Group Reservoir Transformer to predict long-term events more accurately and robustly by overcoming two challenges in Chaos: (1) the extensive historical sequences and (2) the sensitivity to initial conditions. A reservoir is attached to a Transformer to efficiently handle arbitrarily long historical lengths, with an extension of a group of reservoirs to reduce the uncertainty due to the initialization variations. Our architecture consistently outperforms state-of-the-art DNN models in multivariate time series, including NLinear, Pyformer, Informer, Autoformer, and the baseline Transformer, with an error reduction of up to -89.43% in various fields such as ETTh, ETTm, and air quality, demonstrating that an ensemble of butterfly learning, the prediction can be improved to a more adequate and certain one, despite of the traveling time to the unknown future.
1 Introduction
[…]
In the realm of addressing time-series tasks, prior research has explored a multitude of techniques, with a notable emphasis on the utilization of deep neural network models, exemplified by Transformer (Zeng et al., 2022). Nevertheless, an intrinsic constraint associated with the Transformer architecture, originally proposed by (Vaswani et al., 2017), manifests in its quadratic time and memory complexity concerning input length. This inherent limitation imposes a significant hindrance when endeavoring to perform long-term forecasting tasks. Several notable efforts, including UnlimiFormer (Bertsch et al., 2023), Efficient Transformer (Tay et al., 2022), and Reformer (Kitaev et al., 2020), have been made to extend the input length of Transformers. These endeavors primarily hinge on modifications to the attention model itself, often founded on certain assumptions. However, the outcome of these approaches is the extension of the input length to a fixed value, thereby yielding limited adaptability for learning from and predicting arbitrary long sequences.
We directly tackle the challenges associated with stable and accurate long-term chaotic prediction, aiming to push the boundaries of existing technology in time series forecasting. Our approach represents an innovative solution tailored to the complex task of extremely long-term chaotic prediction. We integrate reservoir computing into the Transformer framework, offering a novel approach that can effectively model sequences of inputs with arbitrary lengths. Reservoir computing, known for its simplicity and computational efficiency, excels in processing temporal data within the context of chaotic time series prediction (Bollt, 2021). Within our architecture, the reservoir is instrumental in capturing the intricate dependencies across all temporal events without window-sized context restriction; for instance, it discerns the sequence of events based on all historical data. Subsequently, the Transformer digests short-term event chronicles, melding its insights with those of the reservoir to produce the final forecast. This paradigm departures from conventional heuristic assumptions, ushering in a structured method that propels transformative changes in chaotic forecasting domain.
Our framework grapples with two technical obstacles related to Lyapunov time and the butterfly effect. Firstly, the confined input length of the Transformer considerably constrains the reservoir’s dimensions and potential. This limitation arises since the number of reservoir nodes scales quadratically with the reservoir output size, which mirrors the input size of DNN models, including the Transformer. To efficiently manage longer sequences, our system employs a nonlinear readout rooted in single-head attention mechanisms, superseding the traditional linear reservoir readout. Given its heightened expressiveness, this nonlinear readout essentially performs a dimensionality reduction on reservoir outputs, thus providing the Transformer with more meaningful and potent feature inputs.
[…]
Empirical results demonstrate that the nonlinear readout, when sourced from ensemble reservoirs, drastically refines the Transformer’s prowess in time-series predictions, registering an error reduction of up to 89.43%. Through experiments, we gauge the chaotic nature inherent in our time-series datasets. Furthermore, our methodology surpasses leading-edge DNNs, aptly handling very long input lengths, signaling a transformative phase for the Transformer model.
Our contributions are mainly four folds: a) We integrate reservoir computing into the system, enabling the Transformer to tackle long-term predictions effectively. b) We refine traditional reservoir computing techniques by substituting the linear readout with a nonlinear counterpart, facilitating streamlined dimensionality reduction and adept feature assimilation. c) To address the nuances of reservoir initialization sensitivity, we implement ensemble reservoirs. d) Our approach demonstrated superior performance in forecasting chaotic phenomena, accurately predicting Lyapunov time steps across multiple horizons. Furthermore, it achieved state-of-the-art results in forecasting multivariate time series data.
2 Background
Consider a historical time series data containing $\vec{u}(t)$, a vector of features with $c$ components, and $\vec{y}(t)$ is corresponding labels for each time stamp $t\in T$, where $T$ is the total time stamps in a finite history. Now we aim to forecast $\vec{y}(t+1)$ the class or the value of a future time step given $t$ sequential multivariate history $\vec{u}(1:t+1)$ and $\vec{y}(1:t)$.
Time-Series Forecasting often only handles fixed input length with an assumption of looking back $s$ window size and considering $(t-s)$ history to predict at $t$ instead of using the complete history. Then the task is to predict $\vec{y}(t+1)$ given $\vec{u}(t-s:t+1),\vec{y}(t-s:t)$ with the conditional distribution:
$Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t);\phi)$ (1)
Here $\phi$ indicates learnable parameters shared by all-time steps $T$ to predict the conditional probability. The reason to introduce $s$ is that learning this conditional probability typically depends on $s$. For example, Transformer takes $\mathcal{O}(s^{2}\times c$). The computational cost is quadratic to the $s$.
We introduce deep reservoir computing for time series in chaotic prediction to avoid this costly operation. The advantage of reservoir computing is to preserve all the history for prediction. Therefore, it becomes possible to estimate conditional distribution with any length input:
$Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t);\phi)$ (2)
If the reservoir has $m$ size of the output vector, the time complexity of the proposed work will be $\mathcal{O}((k+m)^{2}\times c)$. Here, the parameter $k$ represents the size of the small look-back window compared to $s$.
2.1 Deep Reservoir Computing
Our basic reservoir architecture follows the deep reservoir computing (DRC) (Gallicchio et al., 2017) with the Echo State Network (ESN) model and a trained linear readout. In the basic Leaky Integrator ESN (LI-ESN) model Gallicchio et al. (2017), the state is updated according to the following state transition function:
$\vec{x}(t)=(1-a)\vec{x}(t-1)+\alpha\tanh(\vec{W}_{in}\vec{u}(t)+\bm{\theta}+\vec{\hat{W}}\vec{x}(t-1)),$ (3)
where $\vec{x}(t)\in R^{n}$ denotes the reservoir state at time $t$, $n$ represents the dimensionality of the reservoir state vector. $\vec{W}_{in}\in R^{c\times n}$ is the input-to-reservoir weight matrix, $\bm{\theta}\in R^{n}$ is the bias-to-reservoir weight vector, $\vec{\hat{W}}\in R^{n\times n}$ is the recurrent reservoir weight matrix, $\tanh(.)$ is the element-wise applied hyperbolic tangent activation function, and $\alpha\in[0,1]$ is the leaky parameter, i.e., the decay rate of the nodes. The reservoir parameters are initialized according to the constraints specified by the Echo State Property (ESP) and then are left untrained Gallicchio & Micheli (2017). Accordingly, the weight values in $\vec{W}_{in}$ and in $\bm{\theta}$ are chosen from a uniform distribution over $[-{scale}_{in},{scale}_{in}]$, where ${scale}_{in}$ represents an input-scaling parameter. Matrix $\vec{\hat{W}}$ contains randomly selected values from a uniform distribution. It is then adjusted so that the spectral radius of matrix $\vec{\tilde{W}}=(1-\alpha)I+\alpha\vec{\hat{W}}$, denoted as $\rho$, remains below 1 ($\rho<1$). This process helps ensure stable dynamics in the reservoir system. The reservoir weight $\vec{\hat{W}}$ and input weight $\vec{W_{in}}$ are fixed and random, where each weight is drawn according to Gaussian distribution with parameterized variances.
2.2 Linear Readout
As reservoir output, a readout component is used to linearly combine the outputs of all the reservoir units as in a standard ESN. The output of a reservoir at each time step $t$ can be computed as
$y(t)=\vec{W}_{out}{\vec{x}(t)}^{T}+\theta_{out},$ (4)
where $y(t)\in R^{m}$ is the linear readout with $m$ dimensionality. $\vec{W}_{out}\in R^{n\times m}$ represents the reservoir-to-readout weight matrix, connecting the reservoir units to the units in the readout. Since only $\vec{W}_{out}$ are trained, the optimization problem boils down to linear regression. Training is typically not a limiting factor in DRC, in sharp contrast with other neural network architectures. The expressiveness and power of reservoir computing rather lie in the high-dimensional non-linear dynamics of the reservoir.
The reservoir is effective for managing lengthy inputs. When we combine it with a Transformer model, the time complexity becomes $\mathcal{O}((k+m)^{2}\times c)$. For the Transformer, if we include more time steps in the input, the training time increases quadratically with the input length. On the other hand, the reservoir treats each time step individually, keeping the entire history in its learning process without making the input longer. This makes the reservoir a better choice for handling long sequences of input data.
2.3 Baseline Transformer
In Transformer, the architecture unravels the temporal dependencies into individual time steps (Wu et al., 2020). At each time of training, we optimize $\phi$ parameters by a Transformer $\mathcal{M}$ to learn the distribution $Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t))$ composed of two inputs: the current time stamp $\vec{u}(t+1)$ and that of the previous $s$ time steps $\vec{u}(t-s:t),\vec{y}(t-s:t)$:
$\vec{g}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)(\vec{u}(t-s:t),\vec{y}(t-s:t))$ (5)
Here $\kappa$ is a learning parameter adjusting the weights of the current input and previous inputs. Thus $\vec{g}(t+1)$ is the input to Transformer, $\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{g}(t+1))$. Nonetheless, employing $\vec{g}(t+1)$ as the input for the Transformer poses two significant challenges. Firstly, the Transformer’s input length is constrained by a factorization approach involving a retrospective window of size $s$, thereby increasing computational complexity. Secondly, the Transformer imposes restrictions on input length, preventing it from exceeding the value of $s$ due to quadratic time complexity considerations. Consequently, the incorporation of extensive historical data becomes impractical, limiting its potential impact on the learning and decision-making processes.
3 Method
3.1 Readout with Self-Attention
One downside of the reservoir is that if the reservoir output size $m$ is large, the Transformer training is slow due to its quadratic complexity. Since a linear readout of a reservoir needs a much larger size to have the same expressive power as the nonlinear readout, we introduce nonlinear readout layers, which work quite well in reducing DRC output dimension and improving prediction performance. Replacing linear readout with nonlinear readout leads to faster convergence and reduces the possibility of getting stuck in a local optimum (Triefenbach & Martens, 2011).
To model nonlinear readout, we use attention mechanisms (Vaswani et al., 2017) to capture input feature importance, alleviating the problem of vanishing gradient of long-distance dependency. There have been various implementations of attention mechanisms. We implement our attention model as in (Zhao, 2022). Attention mechanisms process sequential data that considers the context for each timestamp. $W$ and $b$ are this representation’s weight and bias, and the hyperbolic tangent function $\tanh(.)$ is a nonlinear activation.
$h_{i,j}=\tanh(\vec{x}(t)_{i}^{T}\vec{W_{t}}+\vec{x}(t)_{j}^{T}\vec{W_{x}}+b_{i})$ (6)
$e_{i,j}=\sigma(\vec{W_{\alpha}}h_{i,j}+b_{\alpha})$ (7)
$\gamma_{i,j}=\frac{exp(e_{i,j})}{\sum_{i=1}^{J}(exp(e_{i,j}))}$ (8)
$r(t)=\left\{\sum_{j}\gamma_{i,j}\vec{x}(t)_{j}\right\}_{i=1}^{c}$ (9)
Here, $r(t)$ is a non-linear readout using the attention mechanism and the subscript $i$ and $j$ are the $i$-th and $j$-th components of $\vec{x}(t)$. The attention weight $\gamma$ is a softmax of $e$.
3.2 Group Reservoir
In chaos theory, the butterfly effect is the sensitive dependence on initial conditions in which a small change in one state of a deterministic nonlinear system can result in large differences in a later state. To enhance reservoir performance, we integrate multiple reservoirs’ nonlinear readout layers, ensuring their independence. We consider $Q$ reservoirs with distinct decay rates ($\alpha$ and $\rho$) initialized randomly Gallicchio et al. (2017). Utilizing Equations 3 and 9, if $r^{l}(t)$ represents the nonlinear readout from the attention mechanism of the $l$-th reservoir within $Q$, we combine various leaky parameters ($\alpha$) and spectral radius ($\rho$) across the reservoirs. This results $\vec{o}(t)$ in a grouped ESN representation in element-wise $\oplus$ addition to all reservoirs’ attention outputs:
$\vec{o}(t)=r^{1}(t)\oplus r^{2}(t)\oplus...\oplus r^{Q}(t)$ (10)
3.3 Reservoir Transformer (RT)
Our group reservoir efficiently models observed time-series data in dynamical systems and require minimal training data and computing resources (Gauthier et al., 2021) and is ideal for handling long sequential inputs. Here we mainly describe adopting a reservoir for Transformer Vaswani et al. (2017) as one of the most well-known and valuable DNN architectures. However, our methods are adaptable to other DNN model architectures, for whom we also compared results in our experiments.
Figure a: Non-linear readout for RT.
Figure a: Non-linear readout for RT.
Figure b: Ensemble of multiple reservoirs
Figure 1: Description of deep Reservoir Transformer: In Figure (a), the readout obtained from self-attention, combined with the present input $\vec{u}(t+1)$, serves as the input for training the transformer architecture. In Figure (b), the input $\vec{u}(t)$ is processed individually by $Q$ reservoirs. The nonlinear output of each reservoir is later combined to form the ultimate output, labeled as $\vec{o}(t)$.
Figure 1 shows the architecture of our reservoir-enhanced Transformer. The left figure shows the ensemble of reservoirs, and the right figure shows how to use a non-linear readout reservoir to help Transformer handle arbitrarily long inputs.
Our approach enables comprehensive modeling of entire historical data in decision-making. While the Transformer captures temporal dependencies (Equation 1), limited by a window of size $s$, our method incorporates reservoir pooling to overcome this limitation. This permits the consideration of the entire history without a rise in time complexity. By incorporating Equation 10 (ensemble of all $Q$ reservoir readouts), we arrive at:
$\vec{z}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)\vec{o}(t)$ (11)
$\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{z}(t+1))$ (12)
$\vec{z}(t+1)$ denotes the Transformer $\mathcal{M}$ input with $\kappa$ as a learnable parameter indicating the weights of current multivariate states and reservoir state to predict $\bar{\vec{y}}(t+1)$ by Equation 12.
In the context of multivariate time series prediction, the target label for forecasting up to $q$ future time steps is denoted as $\vec{u}(t+1:t+1+q;v)$, which is contingent upon the observed series $\vec{u}(1:t)$. In the multivariate case, all available features are taken into account with the conditional distribution $Pr(\vec{u}(t+1:t+1+q)|\vec{u}(1:t);\phi)$.
3.4 Training
We enhance the non-linear readout layers through self-attention mechanisms within ensemble reservoirs. This involves updating these layers and performing element-wise addition of the readouts, resulting in an output denoted as $\vec{o}(t)$. This output is concatenated with the present input features $\vec{u}(t+1)$, yielding $z(t+1)$, which then undergoes processing via the Transformer encoder $M(.)$. Initially, we sequentially train the model $M(.)$ across all time stamps $1:T$ using the training dataset. Subsequently, we fine-tune our models using the validation set to achieve optimal model performance. The model is subjected to testing with varying hyper-parameters such as reservoir size, leaky rate, spectral radius, number of reservoirs, learning rate, attention size, Transformer block, and dropout rate. The choice of these hyper-parameters may differ depending on the specific dataset employed (details in the appendix). For a comprehensive understanding of our training algorithm, refer to the pseudo-code provided in the appendix, along with a detailed explanation of the notation employed. In this work, we use the Huber loss function Meyer (2021) as our objective (Equation 13):
$\mathcal{L}(\vec{y},\bar{\vec{y}})=\frac{1}{T}\sum_{i=1}^{T}\begin{cases}\frac{1}{2}(\vec{y}_{i}-\bar{\vec{y}_{i}})^{2},&\text{if}\ |\vec{y}_{i}-\bar{\vec{y}}|\leq\delta\\ \delta|\vec{y}_{i}-\bar{\vec{y}_{i}}|-\frac{1}{2}\delta,&\text{otherwise},\end{cases}$ (13)
Here $\delta$ is an adjustable parameter that controls where the function change occurs to keep the function differentiable, $T$ is the total number of samples and $\vec{y},\bar{\vec{y}}$ are ground truth and predicted values respectively. In the same way, for time series
4 Experiments
We show our experimental results of Reservoir Transformer (RT) compared to baselines, including state-of-the-art methods in time-series prediction, such as: Nlinear (Zeng et al., 2022), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), LogTrans (Li et al., 2019), GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), STMVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017), and Transformer Vaswani et al. (2017). To conduct experiments on multivariate time series, we adhered to the established methodology outlined by  (Zeng et al., 2022), which includes data splitting, preprocessing, and setting horizon times.”
4.1 Result Analysis
We thoroughly assess our RT method on different time series regression tasks. These tasks include Electricity, Traffic, Weather, ETTh1, ETTh2, ETTm1, ETTm2, ILI, Exchange Rate (Zhou et al., 2021), Air Quality (Cini et al., 2021), Daily website visitors (DWV), Daily Gold Price (DGC), and Daily Demand Forecasting Orders (DDFO), Bitcoin Historical Dataset (BTC). We also test RT’s performance on two-time series classification tasks, i.e., Absenteeism at work (AW) and Temperature Readings from IOT Devices (details in the Appendix A.1). In our empirical evaluation, presented in Table 3, a comparison of multivariate long-term forecasting errors using the Mean Squared Error (MSE) metric was performed across various methods and datasets. For each dataset, a forecasting horizon was specified with ILI having horizons $T\in\{24,36,48,60\}$ and the rest having $T\in\{96,192,336,720\}$. It is evident that the RT consistently outperforms other methods, achieving the lowest MSE in most scenarios. The Transformer-based methods also show competitive performance, with some of their results being underscored for distinction (univariate time forecasting experiments in the Appendix A.5).
Test Chaotic Behavior:
Figure a: Mackey Glass Series.
Figure b: ECG Signal
Figure a: Mackey Glass Series.
Figure b: ECG Signal
Figure c: Lorenz Attractor
Figure 2: The figures (a), (b), and (c) represent chaotic datasets. The blue line illustrates the multi-horizon (800) predictions generated by our model, while the green line depicts the actual ground truth data.
Time Series $D_{2}$
Length/Dataset MGS ECG Lorenz MGS ECG Lorenz
50 0.173 0.301 0.313 3.165 2.824 2.425
100 0.185 0.314 0.331 3.242 2.913 2.515
500 0.206 0.341 0.359 3.452 3.173 2.721
1000 0.225 0.374 0.447 3.531 3.282 2.841
Table 1: Comparison of Time Series Lengths and Correlation Dimensions ($D_{2}$) for Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets across varying lengths of 50, 100, 500, and 1000. Higher values of $D_{2}$ indicate increased complexity in the underlying dynamics.
In order to investigate the presence of chaotic behavior and assess the predictive capabilities of our RT models, three distinct chaotic datasets were employed: the Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor. The experiments involved horizons of 50, 100, 500, and 1000, as outlined in Table 1. Our findings indicate that our RT model successfully captures chaotic dynamics across multiple forecasting horizons, shown in Figure 2.
In addition, to show that the datasets we have used are chaotic, we employ the correlation dimension, denoted as $D_{2}$ serves as a reliable measure of chaotic behavior, with higher values reflecting more pronounced chaos. To evaluate the chaotic behavior of various time series datasets  (Pánis et al., 2020), the $D_{2}$ is defined as
$D_{2}=\underset{\epsilon\to 0}{\lim}\frac{\ln{C(\epsilon)}}{\ln{\epsilon}}.$ (14)
The correlation sum $C(\epsilon)$ for some small scalar $\epsilon$ is defined as in (Grassberger & Procaccia, 1983):
$C(\epsilon)=\underset{N\to\infty}{\lim}\frac{2}{N(N-1)}\sum_{i<j}H(\epsilon-|x_{i}-x_{j}|),$ (15)
where $H$ is the Heaviside step function, $N$ is the number of points, and $|x_{i}-x_{j}|$ is the distance between two points . We compute the correlation dimension for all datasets in our experiments using the implementation provided by (Vasquez-Correa, 2015)
Length/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Exchange ILI BTC DGP DWV DDFO
100 0.790 1.190 0.560 0.299 0.995 0.995 0.961 1.070 1.343 0.540
500 1.340 0.973 0.931 0.535 1.099 1.249 0.853 1.626 1.823 0.540
1000 1.266 0.866 1.061 0.667 1.359 1.299 0.853 1.398 1.838 0.540
Table 2: This study investigates the correlation dimension $D_{2}$ across multiple datasets using diverse state numbers (100, 1000, and 10000). The correlation dimension $D_{2}$ serves as an indicator of chaotic behavior, with higher values denoting a more pronounced presence of chaos (Panis, 2020). Our findings indicate that the correlation dimension $D_{2}$ was evaluated for all datasets under varying state numbers.
Table  1 and  2 display the outcomes of a correlation dimension ($D_{2}$) analysis across various datasets. The correlation dimension gauges chaotic tendencies, with higher values suggesting heightened chaos. This assessment involves different state numbers: 50, 100, 500, and 1000 and it exhibits $D_{2}$ values for each dataset and state number. The findings reveal that, for all datasets, the correlation dimension increases with higher state numbers, indicating intensified chaotic behavior. Since Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets are highly chaotic datasets, they showed the higher $D_{2}$.
Model
Horizons
Linear
NLinear
DLinear
FEDformer
Autoformer
Informer
Pyraformer
LogTrans
Repeat
RT
Dataset Electricity 96 0.140 0.141 0.140 0.193 0.201 0.274 0.386 0.258 1.588 0.140
192 0.153 0.154 0.153 0.201 0.222 0.296 0.386 0.266 1.595 0.151
336 0.169 0.171 0.169 0.214 0.231 0.300 0.378 0.280 1.617 0.168
720 0.203 0.210 0.203 0.246 0.254 0.373 0.376 0.283 1.647 0.201
Exchange 96 0.082 0.089 0.081 0.148 0.197 0.847 0.376 0.968 0.081 0.101
192 0.167 0.180 0.157 0.271 0.300 1.204 1.748 1.040 0.167 0.192
336 0.328 0.331 0.305 0.460 0.509 1.672 1.874 1.659 0.305 0.305
720 0.964 1.033 0.643 1.195 1.447 2.478 1.943 1.941 0.823 0.586
Traffic 96 0.410 0.410 0.410 0.587 0.613 0.719 2.085 0.684 2.723 0.391
192 0.423 0.423 0.423 0.604 0.616 0.696 0.867 0.685 2.756 0.399
336 0.436 0.435 0.436 0.621 0.622 0.777 0.869 0.734 2.791 0.425
720 0.466 0.464 0.466 0.626 0.660 0.864 0.881 0.717 2.811 0.442
Weather 96 0.176 0.182 0.176 0.217 0.266 0.300 0.896 0.458 0.259 0.173
192 0.218 0.225 0.220 0.276 0.307 0.598 0.622 0.658 0.309 0.217
336 0.262 0.271 0.265 0.339 0.359 0.578 0.739 0.797 0.377 0.259
720 0.326 0.338 0.323 0.403 0.419 1.059 1.004 0.869 0.465 0.314
ILI 24 1.947 1.683 2.215 3.228 3.483 5.764 1.420 4.480 6.587 0.109
36 2.182 1.703 1.963 2.670 3.103 4.755 7.394 4.799 7.130 0.109
48 2.256 1.719 2.130 2.622 2.669 4.763 7.551 4.800 6.575 0.111
60 2.390 1.819 2.368 2.857 2.770 5.264 7.662 5.278 5.893 0.113
ETTh1 96 0.375 0.374 0.375 0.376 0.449 0.865 0.664 0.878 1.295 0.375
192 0.418 0.408 0.405 0.420 0.500 1.008 0.790 1.037 1.325 0.411
336 0.479 0.429 0.439 0.459 0.521 1.107 0.891 1.238 1.323 0.432
720 0.624 0.440 0.472 0.506 0.514 1.181 0.963 1.135 1.339 0.436
ETTh2 96 0.288 0.277 0.289 0.346 0.358 3.755 0.645 2.116 0.432 0.228
192 0.377 0.344 0.383 0.429 0.456 5.602 0.788 4.315 0.534 0.269
336 0.452 0.357 0.448 0.496 0.482 4.721 0.907 1.124 0.591 0.289
720 0.698 0.394 0.605 0.463 0.515 3.647 0.963 3.188 0.588 0.306
ETTm1 96 0.308 0.306 0.299 0.379 0.505 0.672 0.543 0.600 1.214 0.312
192 0.340 0.349 0.335 0.426 0.553 0.795 0.557 0.837 1.261 0.348
336 0.376 0.375 0.369 0.445 0.621 1.212 0.754 1.124 1.283 0.372
720 0.440 0.433 0.425 0.543 0.671 1.166 0.908 1.153 1.319 0.424
ETTm2 96 0.168 0.167 0.167 0.203 0.255 0.365 0.435 0.768 0.266 0.151
192 0.232 0.221 0.224 0.269 0.281 0.533 0.730 0.989 0.340 0.187
336 0.320 0.274 0.281 0.325 0.339 1.363 1.201 1.334 0.412 0.210
720 0.413 0.368 0.397 0.421 0.433 3.379 3.625 3.048 0.521 0.285
Table 3: Comparing multivariate long-term forecasting errors using MSE: lower values indicate better performance. Specifically, for the ILI dataset, forecasting horizons are $T\in\{24,36,48,60\}$, while for the remaining datasets, they are $T\in\{96,192,336,720\}$. Optimal results are emphasized in bold, while the top results from Transformers are underlined for distinction.”
Horizons 100 500 1500 2500
Dataset Model MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 NLinear RT 0.362 0.371 0.399 0.403 0.439 0.438 0.479 0.442 0.692 0.602 0.798 0.714 1.683 1.294 2.832 2.104
ETTh2 NLinear RT 0.283 0.223 0.350 0.310 0.684 0.303 0.591 0.574 0.947 0.561 1.822 0.935 1.361 0.951 2.891 1.821
ETTm1 NLinear RT 0.302 0.311 0.350 0.352 0.441 0.357 0.433 0.418 0.849 0.739 1.476 1.102 1.153 0.896 2.768 1.562
ETTm2 NLinear RT 0.164 0.127 0.240 0.217 0.411 0.267 0.432 0.329 1.084 0.740 2.148 0.977 2.099 1.180 3.260 1.936
Table 4: Comparison of NLinear and RT for explaining the Lyapunov Exponent’s predictive performance across different datasets. The table presents MSE and MAE values for both models at various prediction horizons: 100, 500, 1500, and 2500. The results demonstrate that the RT model consistently outperforms the NLinear model, achieving significantly lower MSE and MAE values, thus overcoming the Lyapunov Exponent’s challenges of far-sighted forecasting.
Figure 3: Visualization comparing the prediction of the Lyapunov Exponent using NLinear and RT models. The X-axis corresponds to the time sequence, while the Y-axis corresponds to the target feature.
4.2 Ablation Study
Lyapunov Exponent: The Lyapunov Exponent (LE) measures how chaotic a system is by computing how quickly points on a path move away from each other over time. Table 4 shows our RT and a baseline NLinear LE over different time steps of 100, 500, 1500, and 2500, in future on different datasets (ETTh1, ETTh2, ETTm1, ETTm2) and outperforms the NLinear model in predicting the LE. Figure 3 illustrates the comparison of forecasting 2500 future time steps between the NLinear and RT models, where RT matches more the ground truth than NLinear.
Entropy Datasets $\vec{u}(t)$ $\vec{o}(t)$ $\vec{z}(t)$ ETTh1 1.011 1.966 2.331 ETTh2 0.993 1.892 2.015 ETTm1 0.816 2.056 2.831 ETTm2 0.599 1.248 2.131 [Uncaptioned image] Table 5: Top: Entropy in different datasets. This table presents the entropy values of the original inputs ($\vec{u}(t)$), the corresponding reservoir readouts ($\vec{o}(t)$), and the resultant transformed representations ($\vec{z}(t)$) across various datasets. Bottom: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown for ETTh1
[Uncaptioned image] Figure 4: In the presented graph, the X-axis represents the size of the model parameters in thousands, while the Y-axis illustrates the mean absolute error of the prediction outputs. The yellow boxes in the graph indicate the history lengths used in the computation of the RT-1R, RT-5R, and RT-10R models, which were each constructed with one, five, and ten reservoirs, respectively.
Explain Signature Features: We illustrate and measure how well-distributed the features input of Transformer with or without reservoir are. More distribution means higher entropy and more informative, and vice versa. We use the Local Interpretable Model-agnostic Explanations (LIME) algorithm (Ribeiro et al., 2016). Our goal is to explain the decision-making process within the readout layer by employing LIME to generate a local linear approximation of the model. The initial step involves generating perturbed input time series by adding noise to the original $\vec{u(t)}$. A simpler interpretable model is then trained using these perturbed inputs along with corresponding output time series $\vec{y}(t)$. The complexity of the interpretable model depends on the reservoir model’s intricacy, encompassing both linear and non-linear options. Ultimately, we utilize the coefficients of the interpretable model, along with LIME-calculated feature importances, to explain the readout layer’s decision. This Table (Top) 5 shows that the readout from the reservoir, $\vec{o}(t)$, has higher entropy than Transformer original input features $\vec{u}(t)$. By concatenating both using Equation 11, we achieve the highest entropy. Additionally, Table (Bottom) 5 illustrates the Lime explanation for both RT and the baseline Transformer (more ablation in the Appendix A.4).
History length versus parameter size: Conventional approaches for modeling temporal data require a large number of training parameters, proportional to the window size (s) and the number of features (C). This becomes infeasible when considering long historical sequences. In contrast, reservoir computing approaches, such as the proposed method, only train a vector representation (i.e. readout) for the next time step without explicit conditioning on the entire history. Empirical results, illustrated in Figure 4, demonstrate that RT provides a better trade-off between the number of training parameters and loss.
Impact of Reservoir Initial Conditions
Figure 5: Comparison of MSE loss across five reservoir initialization techniques: random initialization, zero initialization, constant initialization, normal distribution initialization, and uniform distribution initialization. The x-axis represents the initialization methods, while the y-axis denotes the corresponding MSE loss values.
In this study, we delved into the intricate dynamics of reservoir initialization and its consequential impact on the performance of our RT models. The reservoir initialization process lays the foundation for the system’s dynamic behavior, influencing the trajectory of internal states and subsequent information processing capabilities. We investigated the impact of different initialization conditions on the performance of our RT model. We employed five distinct initialization approaches, both in ensemble configurations with group reservoirs and without group reservoirs, as depicted in Figure 5.
Firstly, we utilized random initialization, which initializes the reservoir randomly within the range of 0 to 1. Subsequently, we employed zero initialization, which sets all reservoir values to zero. We also explored constant initialization, where a constant value is randomly selected for initialization. Furthermore, we investigated initialization using normal and uniform distributions.
Our experimentation revealed that among these initialization methods, zero initialization consistently yielded superior results in terms of MSE loss, both in ensemble setups with and without group reservoirs. Notably, the ensemble technique with group reservoirs demonstrated the lowest loss across both the ETTh2 and ETTm2 datasets.
5 Related Work
Chaotic time-series prediction using machine learning approaches has been investigated in the last decade using CNN, WNN, FNN, LSTM, etc. (Ramadevi & Bingi, 2022). For example,(Sangiorgio & Dercole, 2020) proposed the use of a Recurrent Neural Network (RNN) with Long Short-Term Memory (LSTM) cells for chaotic time-series prediction and demonstrated improved performance on several chaotic systems.(Karasu & Altan, 2022) showed a new type of LSTM cell to handle chaotic time series prediction and reported that it outperforms traditional LSTM on several chaotic systems. Despite the domain knowledge studied in broad areas, developing new machine learning technologies for Chao’s purpose still needs more attention. In particular, the connection between chaotic prediction and conventional time-series forecasting is yet to be more closely integrated (Sahoo et al., 2019). There has been extensive literature on time-series forecasting as well  (PaperwithCodes, 2022), and sometimes these datasets hold chaotic properties as well, including using Transformer- and non-Transformer-based solutions for long-term time series forecasting (LTSF). For example,  (Mohammadi Farsani & Pazouki, 2020) showed the use of self-attention in the Transformer model for time-series prediction tasks that improved performance. Similarly,  (Huang et al., 2019) proposed the use of self-attention in a Recurrent Neural Network (RNN) architecture for time-series prediction, achieving improved performance on several datasets. However, these approaches still need more effective models for long-sequential input handling to consider the long history of the past. We have proposed a reservoir Transformer (RT) to reach this aim.
While reservoirs have found extensive application in time-series tasks, their performance often lags behind the current state-of-the-art outcomes (Gauthier, 2021). A concept of Reservoir Transformer  (Shen et al., 2020) was introduced by interspersing the reservoir layers into the regular Transformer layers and showing improvements in wall-clock compute time. However, such integration will reduce the feature extraction power of both reservoirs in the long-sequence and Transformer in short sequences. We do not combine the reservoir and Transformer linearly but in a non-linear way and receive significant improvements over all previous methods in time-series prediction, including chaotic prediction.
Additionally, there has been research on using ensembles of models for time-series prediction.(Feng et al., 2019) proposed an ensemble method that combines multiple hybrid models using wavelet transform for chaotic time-series prediction and achieved improved performance on several chaotic systems.(Tang et al., 2020) proposed an ensemble method that combines multiple deep-learning models for time-series prediction and showed improved performance on several datasets. In our work, we apply ensemble techniques to our Reservoir Transformer for the first time, contributing to further advancements in time-series prediction.
6 Conclusion
We introduce nonlinear readout to reservoir computing that significantly enhances the performance of the state-of-the-art DNNs, including Transformer. Our deep Reservoir Transformer has the unique advantage of handling any input length and producing robust prediction outputs. Thus, it is ideal for chaotic time-series forecasting. This research will open new research questions on seamlessly integrating chaotic analysis, DNNs, and reservoir computing to better understand and predict our environments.
Appendix A Appendix
A.1 Dataset and Parameter Configuration
In this subsection, we outline the approach taken for dataset partitioning and parameter selection within our research framework.
Dataset Partitioning: The dataset is divided sequentially into three distinct sets: a training set, a validation set, and a test set. Specifically, the first $70\%$ of the data is assigned to the training set, while $10\%$ is allocated to the validation set, and the remaining $20\%$ forms the test set. This partitioning strategy enables effective model training, validation, and evaluation.
Model Parameters: The configuration of model parameters plays a crucial role in our research methodology. For the Transformer architecture, we employ four Transformer blocks, each comprising 32 multi-head attention mechanisms with a total of four attentions. In the case of the Feedforward Neural Network (FNN), we utilize hidden layers containing 64 units each.
Within the network architecture, dropout is employed as a regularization technique. Specifically, a dropout rate of 0.4 is applied to the hidden layers, while a rate of 0.1 is used for the reservoir network. Additionally, dropout with a rate of 0.2 is integrated into the attention network.
Furthermore, the reservoir state size varies between 20 and 50 units across different ensembling reservoirs. To introduce diversity in the ensembling process, sparsity levels range from 0.3 to 0.7 for different reservoirs, accompanied by leaky units with values ranging from 0.2 to 0.6.
Finally, a learning rate of $1\times 10^{-4}$ is adopted to facilitate the training process and achieve optimal convergence.
By meticulously configuring these parameters, we aim to attain a comprehensive understanding of the model’s performance across various experimental setups.
Evaluation metric: In order to analyze the performance and evaluation of the proposed work, Mean Squared Error (MSE), and Mean Absolute Error (MAE) has been used for regression. The MSE is a quantification of the mean squared difference between the actual and predicted values. The minimum difference, the better the model’s performance. On the other hand, MAE is a quantification of the mean difference between the actual and predicted values. As like as MSE, the minimum score of MAE, the better the model’s performance. For classification tasks, we have used Accuracy and F1 score metrics.
Dataset Description:
  • ETTh (Electricity Transformer Temperature hourly-level): The ETTh dataset presents a comprehensive collection of transformer temperature and load data, recorded at an hourly granularity, covering the period from July 2016 to July 2018. The dataset has been meticulously organized to encompass hourly measurements of seven vital attributes related to oil and load characteristics for electricity transformers. These incorporated features provide a valuable resource for researchers to glean insights into the temporal behavior of transformers.
    For the ETTh1 subset of the dataset, our model architecture entails the utilization of reservoir units ranging in size from 20 to 50, coupled with an ensemble comprising five distinct reservoirs. The attention mechanism deployed in this context employs a configuration of 15 attention heads, each possessing a dimension of 10 for both query and key components.
    Likewise, for the ETTh2 subset, our approach involves employing a total of seven ensemble reservoirs, each incorporating reservoir units spanning from 20 to 50 in size. The remaining parameters for this subset remain consistent with the configuration established for ETTh1.
  • ETTm (Electricity Transformer Temperature 15-minute-level): The ETTm dataset offers a heightened level of detail and granularity by capturing measurements at intervals as fine as 15 minutes. Similar to the ETTh dataset, ETTm spans the timeframe from July 2016 to July 2018. This dataset retains the same seven pivotal oil and load attributes for electricity transformers as in the ETTh dataset; however, the key distinction lies in its higher temporal resolution. This heightened level of temporal granularity holds the potential to unveil subtler patterns and intricacies in the behavior of transformers.
    In the context of model architecture, the configuration adopted for the ETTm dataset involves the use of ensemble reservoirs. Specifically, for the ETTm1 subset, an ensemble of seven reservoirs is employed, while for the ETTm2 subset, an ensemble of six reservoirs is utilized. This distinctive architectural setup is designed to harness the increased temporal resolution of the ETTm dataset, enabling the model to capture and interpret the finer nuances within the transformer data.
  • Exchange-Rate This dataset includes the daily exchange rates for eight foreign countries: Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore. This data ranges from the year 1990 to 2016. To analyze this information, we are using 15 reservoirs, which are a specific type of model to help us understand and predict changes in these exchange rates.
  • AQI (Air Quality ): The AQI dataset encompasses a collection of recordings pertaining to various air quality indices, sourced from 437 monitoring stations strategically positioned across 43 cities in China. Specifically, our analysis focuses solely on the PM2.5 pollutant for this study. Notably, previous research efforts in the realm of imputation have concentrated on a truncated rendition of this dataset, which comprises data from 36 sensors.
    For our investigation, we leverage 15 ensemble reservoirs as the architectural foundation for this dataset. This ensemble configuration is tailored to the distinctive characteristics and complexities inherent in the AQI data. By deploying this setup, we endeavor to enhance our capacity to effectively address the intricate imputation challenges associated with the AQI dataset.
  • Daily Website visitors: [^1] https://www.kaggle.com/datasets/bobnau/daily-website-visitorsThe dataset pertinent to daily website visitors provides a comprehensive record of time series data spanning five years. This dataset encapsulates diverse metrics of traffic associated with statistical forecasting teaching notes. The variables encompass daily counts encompassing page loads, unique visitors, first-time visitors, and returning visitors to an academic teaching notes website. The dataset comprises 2167 rows, covering the timeline from September 14, 2014, to August 19, 2020.
    From this array of variables, our specific focus centers on the ’first-time visitors,’ which we have identified as the target variable within the context of a regression problem. By delving into this specific variable, we aim to uncover insights regarding the behavior of new visitors to the website and its associated dynamics.
    In terms of the model architecture employed for this dataset, we have constructed a configuration centered around 12 ensemble reservoirs. This architectural choice reflects the dataset’s intricacies and the requirement for an approach that can effectively capture the underlying patterns and variations within the daily website visitor data
  • Daily Gold Price:  [^2] https://www.kaggle.com/datasets/nisargchodavadiya/daily-gold-price-20152021-time-series The dataset pertaining to daily gold prices offers an extensive collection of time series data encompassing a span of five years. This dataset encapsulates a range of metrics pertaining to gold price dynamics. The variables included cover daily measurements related to various aspects of gold pricing. The dataset consists of 2167 rows, covering the time period from September 14, 2014, to August 19, 2020.
    The primary objective in this analysis is regression, wherein we seek to model relationships and predict outcomes based on the available variables. Specifically, we are interested in predicting the daily gold price based on the given set of features.
    In terms of the architectural setup for this dataset, we have established a model configuration that employs 10 ensemble reservoirs. This choice is rooted in the need to effectively capture the nuances and trends inherent in the daily gold price data. By employing this ensemble-based approach, we aim to enhance the model’s capacity to discern patterns and fluctuations within the gold price time series.
  • Daily Demand Forecasting Orders:  [^3] https://archive.ics.uci.edu/dataset/409/dailydemandforecastingorders This is a regression dataset that was collected during 60 days, this is a real database of a Brazilian company of large logistics. Twelve predictive attributes and a target that is considered a non-urgent order. We used 17 ensemble reservoirs for this dataset.
  • Bitcoin Historical Dataset (BTC):  [^4] https://finance.yahoo.com/quote/BTC-USD/history/ The dataset used in this research contains the historical price data of Bitcoin from its inception in 2009 to the present. It includes key metrics such as daily opening and closing prices, trading volumes, and market capitalizations. This comprehensive dataset captures Bitcoin’s significant fluctuations and trends, reflecting major milestones and market influences. It serves as an essential tool for analyzing patterns, understanding price behaviors, and predicting future movements in the cryptocurrency market. We used 10 ensemble reservoirs for this dataset.
  • Absenteeism at Work:  [^5] https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work The dataset focusing on absenteeism at work is sourced from the UCI repository, specifically the ”Absenteeism at work” dataset. In our analysis, the dataset has been utilized for classification purposes, specifically targeting the classification of individuals as social drinkers. The dataset comprises 21 attributes and encompasses a total of 740 records. These records pertain to instances where individuals were absent from work, and the data collection spans the period from January 2008 to December 2016.
    Within the scope of this analysis, the primary objective is classification, wherein we endeavor to categorize individuals as either social drinkers or non-social drinkers based on the provided attributes.
    The architectural arrangement for this dataset involves the utilization of 15 ensemble reservoirs. This choice of model configuration is underpinned by the aim to effectively capture the intricacies inherent in the absenteeism data and its relationship to social drinking behavior. Employing ensemble reservoirs empowers the model to discern nuanced patterns and interactions among the attributes, thereby enhancing its classification accuracy.
  • Temperature Readings: IOT Devices:  [^6] https://www.kaggle.com/datasets/atulanandjha/temperature-readings-iot-devices The dataset pertaining to temperature readings from IoT devices centers on the recordings captured by these devices, both inside and outside an undisclosed room (referred to as the admin room). This dataset comprises temperature readings and is characterized by five distinct features. In total, the dataset encompasses 97605 samples.
    The primary objective in this analysis involves binary classification, where the aim is to classify temperature readings based on whether they originate from an IoT device installed inside the room or outside it. The target column holds the binary class labels indicating the origin of the temperature reading.
    In terms of dataset characteristics, the architecture of our model involves the utilization of 15 ensemble reservoirs. This architectural choice is rooted in the desire to effectively capture the underlying patterns and distinctions present within the temperature readings dataset. The use of ensemble reservoirs enhances the model’s capacity to distinguish between the different temperature reading sources, thereby facilitating accurate classification.
A.2 Pseudo Algorithm For Deep Reservoir Transformer
Algorithm 1 Training algorithm of Deep Reservoir Computing
$\vec{u}(1:T,1:c),\vec{y}(1:T)$
Optimize $Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t))$ distribution by learning a model $M(.)$
$\vec{W}_{in},\hat{\vec{W}}\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir
$\sigma\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir non-linear output
$\phi\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Transformer encoder $M(.)$
while $epoch<epochs$ do
     for $i\leftarrow 1$ to $T$ do
         for $k\leftarrow 1$ to $GroupESN$ do
              $o^{k}(i)\leftarrow R_{k}(\vec{u}(i),\vec{y}(i);\vec{W}_{in},\hat{\vec{W}},\sigma)$triangleright Non-linear readout from $k-th$ reservoir
              $o(i)\leftarrow o(i)\oplus o^{k}(i)$triangleright Element-wise addition among all reservoirs
         end for
         $\vec{z}(i+1)\leftarrow\kappa\vec{u}(i+1)+(1-\kappa)o(i+1)$triangleright Concatenation
         $\bar{\vec{y}}(i+1)\leftarrow\mathcal{M}(\vec{z}(i+1);\phi)$
     end for
     $loss\leftarrow\mathcal{L}(\bar{\vec{y}}(i+1),\vec{y}(i+1))$triangleright Calculating loss for every time step $z$
     $\phi,\sigma\leftarrow update$triangleright Update parameters
end while
The proposed algorithm 1 is for training a deep reservoir computing model, with the goal of optimizing the distribution of the target variable, $\vec{y}(t+1)$, given the input variables, $\vec{u}(1:t+1)$, and previous target variables, $\vec{y}(1:t)$. The algorithm initializes the weights of the reservoir, $\vec{W}_{in}$ and $\hat{W}$, as well as the weights of the non-linear output, $\sigma$, and the Transformer encoder, $\phi$, as random variables sampled from a normal distribution with mean 0 and standard deviation 1.
The algorithm then enters an outer loop, where the number of iterations is determined by the number of epochs. Within this loop, there is an inner loop that iterates through each time step, $i$, from 1 to $T$. For each time step, a group of reservoirs, $GroupESN$, are processed in parallel. Each reservoir, $R_{k}$, receives the input, $\vec{u}(i)$, and the previous target variable, $\vec{y}(i)$, as input, and calculates a non-linear readout, $o^{k}(i)$, using the weights $\vec{W}_{in},\hat{\vec{W}}$, and $\sigma$. The readouts from all reservoirs are then element-wise added together to produce $o(i)$.
The input of the next time step, $\vec{u}(i+1)$, is concatenated with the readout from the reservoirs, $o(i)$, using a parameter $\kappa$, to produce $\vec{z}(i+1)$. Finally, the Transformer encoder, $\mathcal{M}(.)$, is applied to $\vec{z}(i+1)$ using the parameters $\phi$, to produce the target variable for the next time step, $\vec{y}(1+1)$.
The loss is calculated for every time step, $z$, by comparing the predicted target variable, $\bar{\vec{y}}(1:T)$, with the true target variable, $\vec{y}(1:T)$, using a loss function, $\mathcal{L}(.)$. The parameters, $\phi$ and $\sigma$, are then updated to minimize the loss.
A.3 Variable Descriptions
Variable Description
$\mathcal{M}(.)$
A Transformer or DNN model.
$\vec{u}$
Samples with $c$ components.
$c$
Number of features in dataset
$\vec{y}$
Target features for prediction.
$\bar{\vec{y}}$
Prediction value
$T$
Entire time span or duration
$t$
Represents a specific moment or data point in time
$R$
Non-linear readout representation of reservoir
$\vec{o}$
Aggregate output from all reservoirs within the group
$\rho$
Spectral radius of reservoir
$Q$
Number of reservoirs in GroupESN
$m$
Reservoir output size
$\vec{W}_{i}n$
Input-to-reservoir weight matrix
${\vec{W}_{out}}$
Reservoir to readout weight matrix
${\theta_{out}}$
Reservoir to readout bias matrix
$\vec{\hat{W}}$
Recurrent reservoir weight matrix
$\vec{x}$
Reservoir state
$\kappa$
Parameter to indicate the weights of current state and reservoir state
$\phi$
Learnable parameters of $M(.)$
$s$
Window size to look back previous sequence
$R(.)$
Deep reservoir computing
$d$
Feature dimension of Transformer
$\vec{g}$
Input for the baseline Transformer
$\vec{z}$
The concatenation of reservoir and current multivariate
$\mathcal{L}$
Loss function
$D_{2}$
Correlation dimension
$C(\epsilon)$
Correlation sum
$H(.)$
Heaviside step function
Table 6: Descriptions of all variables used in this paper
A.4 Ablation Study
Figure 6: The graph illustrates the relationship between the number of ensemble reservoirs used and the mean square error (MSE) of the prediction outputs. The X-axis represents the number of reservoirs included in the ensemble, while the Y-axis depicts the corresponding MSE values. The ’star’ in the graph denotes the minimum MSE achieved.
A.4.1 Impact of Reservoir Count on Performance:
Figure 6 illustrates the relationship between the number of ensembling reservoirs and system performance. The x-axis depicts the count of reservoirs, while the y-axis portrays the Mean Squared Error (MSE) value. The graph discernibly demonstrates that increasing the number of reservoirs positively correlates with performance enhancement, leading to a reduction in loss. However, beyond a certain threshold of reservoir count (best hyperparameters of the number of reservoirs), the performance improvement saturates, resulting in a plateau in loss reduction. This phenomenon suggests an optimal range for reservoir count in relation to performance enhancement.
A.4.2 Graphical Explanation of Features
In Figure 7, we present a visual depiction that offers a comprehensive insight into the interpretability of features derived from both the reservoir output $\vec{o}(t)$ and input feature $\vec{u}(t)$.
Figure 7: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown. The test dataset is represented by ETTh1, ETTh2, ETTm1, and ETTm2. The results indicate that current input has a greater impact on feature 6, while the reservoir affects multiple features such as 20, 13, 29, and 7.
A.5 Experiments on Univariate Time sSeries
Table 7 shows that RT model outperforms GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), ST-MVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017) on the air quality dataset. On regression and classification tasks, RT consistently outperforms the baseline Transformer with respect to all criteria, achieving significantly lower MSE and MAE values across various datasets like DWV, DGP, DDFO, and BTC. This suggests that the RT model is better equipped to capture complex temporal patterns. In the classification tasks, the RT model demonstrates better performance in terms of accuracy, F1 Score, Jaccard Score, Roc Auc, Precision, and Recall across different datasets like Absenteeism at work and Temperature Readings. Through this analysis, we provide evidence that the RT architecture excels not only in regression tasks but also in classification tasks, outperforming traditional Transformer architectures.
Air Quality
Metric/Model GRIN BRITS ST-MVL M-RNN ImputeTS RT
MAE 10.514 12.242 12.124 14.242 19.584 9.046
Method Training Test
REGRESSION TASKS
Model/Metric MSE MAE R2 MSE MAE R2 Parameters
Daily website visitors (DWV)
Baseline 98.451 7.583 0.996 42.954 5.181 0.997 $\sim$13.2k
RT 10.400 (-89.43%) 2.114 0.999 6.181 1.893 0.999 $\sim$3.2k
Daily Gold Price (DGP)
Baseline 39459.790 112.251 0.996 168608.260 353.310 0.959 $\sim$13.9k
RT 22022.981 86.985 0.998 28790.317 119.227 0.994 $\sim$3.4k
Daily Demand Forecasting Orders (DDFO)
Baseline 312.933 13.274 0.929 1013.962 22.265 0.099 $\sim$14.1k
RT 76.385 7.487 0.985 410.652 10.173 0.720 $\sim$3.7k
Bitcoin Historical Dataset (BTC)
Baseline 188614.931 427.801 0.968 79389.491 232.119 0.961 $\sim$10.5k
RT 158651.081 310.918 0.998 74800.535 211.626 0.989 $\sim$2.8k
CLASSIFICATION TASKS:
Model Accuracy F1 Score Jaccard Score Roc Auc Precision Recall Parameters
Absenteeism at work
Baseline 0.928 0.940 0.887 0.929 0.953 0.927 $\sim$14.2k Training
RT 0.947 0.956 0.916 0.950 0.976 0.936 $\sim$3.5k Training
Baseline 0.844 0.839 0.722 0.844 0.833 0.845 $\sim$14.2k Test
RT 0.899 0.901 0.819 0.901 0.944 0.860 $\sim$3.5k Test
Temperature Readings
Baseline 0.888 0.934 0.877 0.877 0.877 0.915 $\sim$10.9k Training
RT 0.959 0.975 0.952 0.934 0.981 0.970 $\sim$2.3k Training
Baseline 0.649 0.688 0.524 0.646 0.632 0.755 $\sim$10.9k Test
RT 0.802 0.843 0.728 0.794 0.867 0.820 $\sim$2.3k Test
Table 7: Assessing the performance of regression on both training and test datasets, we utilize a window size of $s=5$ to observe historical data in baseline Transformers.

### v2

@@
-Md Kowsher School of Engineering and Science
+Changes by Butterflies: Farsighted Forecasting with Group Reservoir Transformer
 Abstract
@@
 We introduce Group Reservoir Transformer to predict long-term events more accurately and robustly by overcoming two challenges in Chaos: (1) the extensive historical sequences and (2) the sensitivity to initial conditions.
-A reservoir is attached to a Transformer to efficiently handle arbitrarily long historical lengths, with an extension of a group of reservoirs to reduce the uncertainty due to the initialization variations.
-Our architecture consistently outperforms state-of-the-art DNN models in multivariate time series, including NLinear, Pyformer, Informer, Autoformer, and the baseline Transformer, with an error reduction of up to -89.43% in various fields such as ETTh, ETTm, and air quality, demonstrating that an ensemble of butterfly learning, the prediction can be improved to a more adequate and certain one, despite of the traveling time to the unknown future.
+A reservoir is attached to a Transformer to efficiently handle arbitrarily long historical lengths, with an extension of a group of reservoirs to reduce the sensitivity to the initialization variations.
+Our architecture consistently outperforms state-of-the-art models in multivariate time series, including TimeLLM, GPT2TS, PatchTST, DLinear, TimeNet, and the baseline Transformer, with an error reduction of up to -59% in various fields such as ETTh, ETTm, and air quality, demonstrating that an ensemble of butterfly learning can improve the adequacy and certainty of event prediction, despite of the traveling time to the unknown future.
 1 Introduction
 […]
-In the realm of addressing time-series tasks, prior research has explored a multitude of techniques, with a notable emphasis on the utilization of deep neural network models, exemplified by Transformer (Zeng et al., 2022).
-Nevertheless, an intrinsic constraint associated with the Transformer architecture, originally proposed by (Vaswani et al., 2017), manifests in its quadratic time and memory complexity concerning input length.
-This inherent limitation imposes a significant hindrance when endeavoring to perform long-term forecasting tasks.
-Several notable efforts, including UnlimiFormer (Bertsch et al., 2023), Efficient Transformer (Tay et al., 2022), and Reformer (Kitaev et al., 2020), have been made to extend the input length of Transformers.
-These endeavors primarily hinge on modifications to the attention model itself, often founded on certain assumptions.
-However, the outcome of these approaches is the extension of the input length to a fixed value, thereby yielding limited adaptability for learning from and predicting arbitrary long sequences.
-We directly tackle the challenges associated with stable and accurate long-term chaotic prediction, aiming to push the boundaries of existing technology in time series forecasting.
-Our approach represents an innovative solution tailored to the complex task of extremely long-term chaotic prediction.
-We integrate reservoir computing into the Transformer framework, offering a novel approach that can effectively model sequences of inputs with arbitrary lengths.
+Prior research has explored a multitude of techniques in chaotic event prediction, with a notable emphasis on the utilization of deep neural network models, exemplified by Transformer (Zeng et al., 2022).
+Nevertheless, the Transformer architecture’s quadratic time and memory complexity of the input length significantly limits the performance of long-term forecasting tasks (Vaswani et al., 2017).
+Several notable efforts, including but not limited to UnlimiFormer (Bertsch et al., 2023), Efficient Transformer (Tay et al., 2022), and Reformer (Kitaev et al., 2020), extend the input length of Transformers but primarily hinge on modifications to the attention model itself, often founded on certain assumptions to extend the input length to a fixed value without considering the temporal dependency of the very far away events, thereby yielding limited adaptability for learning from and predicting arbitrary long sequences.
+Stream reservoir learning: We integrate reservoir computing, as an autoregressive model, into the Transformer framework that effectively models sequences of inputs with arbitrary lengths.
 Reservoir computing, known for its simplicity and computational efficiency, excels in processing temporal data within the context of chaotic time series prediction (Bollt, 2021).
-Within our architecture, the reservoir is instrumental in capturing the intricate dependencies across all temporal events without window-sized context restriction; for instance, it discerns the sequence of events based on all historical data.
-Subsequently, the Transformer digests short-term event chronicles, melding its insights with those of the reservoir to produce the final forecast.
-This paradigm departures from conventional heuristic assumptions, ushering in a structured method that propels transformative changes in chaotic forecasting domain.
-Our framework grapples with two technical obstacles related to Lyapunov time and the butterfly effect.
-Firstly, the confined input length of the Transformer considerably constrains the reservoir’s dimensions and potential.
-This limitation arises since the number of reservoir nodes scales quadratically with the reservoir output size, which mirrors the input size of DNN models, including the Transformer.
-To efficiently manage longer sequences, our system employs a nonlinear readout rooted in single-head attention mechanisms, superseding the traditional linear reservoir readout.
-Given its heightened expressiveness, this nonlinear readout essentially performs a dimensionality reduction on reservoir outputs, thus providing the Transformer with more meaningful and potent feature inputs.
+Within our architecture, the reservoir captures dependencies across all temporal events by discerning the sequence of events based on all historical data without window-sized context restriction.
+The reservoir learning processes the input in a stream fashion and requires no training by freezing reservoir weights.
+Its linear time complexity and the constant space complexity to the input length make handling the arbitrary context length possible.
+Subsequently, the Transformer digests short-term events in quadratic time, melding its insights with long-term events modeled by the reservoir to produce a more accurate final forecast.
+Our framework grapples with two technical obstacles related to Lyapunov time and the butterfly effect using the following two techniques: nonlinear readout and group reservoir.
+Non-linearity: The traditional linear readout is expensive since the number of reservoir nodes scales quadratically with the reservoir output size, which mirrors the input size of the Transformer.
+Our system employs a nonlinear readout rooted in single-head attention mechanisms to manage longer sequences efficiently, superseding the traditional linear reservoir readout.
+Given its heightened expressiveness, this nonlinear readout reduces reservoir outputs’ dimensionality, thus providing the Transformer with more meaningful and potent feature inputs and dimension reduction.
 […]
-Empirical results demonstrate that the nonlinear readout, when sourced from ensemble reservoirs, drastically refines the Transformer’s prowess in time-series predictions, registering an error reduction of up to 89.43%.
-Through experiments, we gauge the chaotic nature inherent in our time-series datasets.
-Furthermore, our methodology surpasses leading-edge DNNs, aptly handling very long input lengths, signaling a transformative phase for the Transformer model.
-Our contributions are mainly four folds: a) We integrate reservoir computing into the system, enabling the Transformer to tackle long-term predictions effectively.
-b) We refine traditional reservoir computing techniques by substituting the linear readout with a nonlinear counterpart, facilitating streamlined dimensionality reduction and adept feature assimilation.
-c) To address the nuances of reservoir initialization sensitivity, we implement ensemble reservoirs.
-d) Our approach demonstrated superior performance in forecasting chaotic phenomena, accurately predicting Lyapunov time steps across multiple horizons.
+Empirical results demonstrate that the nonlinear readout and group reservoir drastically improve the Transformer and its variants in time-series predictions with an error reduction of up to 59.70%.
+Our methodology surpasses Transformer and other leading-edge deep neural networks (DNNs).
+It aptly handles very long input lengths, signaling a notable improvement for the Transformer model.
+Our contributions are mainly fourfold: (1) We integrate reservoir computing into the system, enabling the Transformer to tackle long-term predictions effectively.
+(2) We refine traditional reservoir computing techniques by improving the linear readout with a nonlinear counterpart, facilitating streamlined dimensionality reduction while keep the autoregressive learning on temporal patterns.
+(3) We implement the group reservoir to address the nuances of reservoir initialization sensitivity.
+(4) Our approach demonstrated superior performance in forecasting chaotic phenomena in long-term horizons.
 Furthermore, it achieved state-of-the-art results in forecasting multivariate time series data.
 2 Background
-Consider a historical time series data containing $\vec{u}(t)$, a vector of features with $c$ components, and $\vec{y}(t)$ is corresponding labels for each time stamp $t\in T$, where $T$ is the total time stamps in a finite history.
-Now we aim to forecast $\vec{y}(t+1)$ the class or the value of a future time step given $t$ sequential multivariate history $\vec{u}(1:t+1)$ and $\vec{y}(1:t)$.
-Time-Series Forecasting often only handles fixed input length with an assumption of looking back $s$ window size and considering $(t-s)$ history to predict at $t$ instead of using the complete history.
-Then the task is to predict $\vec{y}(t+1)$ given $\vec{u}(t-s:t+1),\vec{y}(t-s:t)$ with the conditional distribution:
-$Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t);\phi)$ (1)
-Here $\phi$ indicates learnable parameters shared by all-time steps $T$ to predict the conditional probability.
-The reason to introduce $s$ is that learning this conditional probability typically depends on $s$.
-For example, Transformer takes $\mathcal{O}(s^{2}\times c$).
-The computational cost is quadratic to the $s$.
-We introduce deep reservoir computing for time series in chaotic prediction to avoid this costly operation.
-The advantage of reservoir computing is to preserve all the history for prediction.
-Therefore, it becomes possible to estimate conditional distribution with any length input:
-$Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t);\phi)$ (2)
-If the reservoir has $m$ size of the output vector, the time complexity of the proposed work will be $\mathcal{O}((k+m)^{2}\times c)$.
-Here, the parameter $k$ represents the size of the small look-back window compared to $s$.
+Preliminary: We consider a time series dataset denoted by $\mathbf{u}_{1:T}$, where each element $\mathbf{u}_{t}\in\mathbb{R}^{N_{u}}$ represents a $N_{u}$-dimensional vector at time $t\in\{1,\cdots,T\}$.
+Given a rolling forecasting scenario with a fixed window size $w$, our objective is to predict $\tau$ future time steps, $\mathbf{u}_{t+1:t+\tau}$, based on the input $\mathbf{u}_{t-w:t}$.
+We propose the following model:
+$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{u}_{t-w:t},\mathbf{\Phi})$ (1)
+Here, $\mathbf{\Phi}$ represents learnable parameters, and $\mathbb{M}(\cdot)$ is a Transformer time-series forecasting model.
+3 Method
+Transformer has a quadratic time complexity in terms of the history length and prediction future time steps (Jin et al., 2024; Nie et al., 2022; Beltagy et al., 2020).
+To reduce the computational complexity of window sizes, we introduce reservoir computing $\mathcal{R}$ to model the entire history of the time series efficiently:
+$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{u}_{t-k:t}\uplus\mathcal{R}(\mathbf{u}_{1:T}),\mathbf{\Phi})$ (2)
+[…]
+3.1 Cross-attention for Embedding
+We first embed the time series data for the neural networks.
+Since the embedding window has a fixed size, time series sequential input usually are chunked into segments, and the learning happens segment by segment.
+To make each segment comparable to the others, the full time series input are often normalized.
+Here we use the reversible instance normalization technique for normalization (Kim et al., 2021; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021), where the full sequence $\mathbf{u}_{1:T}$ is normalized to have zero mean and unit variance.
+We denote each token embedding as $\epsilon(\mathbf{u_{t}})\in\mathbb{R}^{d_{\epsilon}}$, where $\quad\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$, and the input $\mathbf{u}_{1:T}$ is embedded into ${\epsilon({\mathbf{u}}_{1}),\epsilon({\mathbf{u}}_{2}),\ldots,\epsilon({\mathbf{u}}_{T}})$.
+To address the dependency of each time step in one segment, we apply the cross-attention (Zhang and Yan, 2022; Rombach et al., 2022), where each vector $\epsilon(\mathbf{u}_{t-k:t})$ considers the local temporal relationship with its $i$-window-sized neighbors $\epsilon(\mathbf{u}_{t-i:t+i})$:
+$\mathbf{h}_{t}=\text{softmax}\left(\frac{\mathbf{W}_{k}\cdot\epsilon(\mathbf{u}_{t-k:t})(\mathbf{W}_{q}\cdot\epsilon(\mathbf{u}_{t-i:t+i}))^{T}}{\sqrt{d_{\epsilon}}}\right)\mathbf{W}_{v}\cdot\epsilon(\mathbf{u}_{t-i:t+i}),$ (3)
+where $\mathbf{h}_{t}\in\mathbb{R}^{k\times d_{\epsilon}}$ carries the local temporal relationship for time step $t$.
+Here $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, $\mathbf{W}_{v}$ are trainable matrices to learn queries, keys, and values, respectively (Wei et al., 2020).
 3.2 Deep Reservoir Computing
-Our basic reservoir architecture follows the deep reservoir computing (DRC) (Gallicchio et al., 2017) with the Echo State Network (ESN) model and a trained linear readout.
-In the basic Leaky Integrator ESN (LI-ESN) model Gallicchio et al.
-(2017), the state is updated according to the following state transition function:
-$\vec{x}(t)=(1-a)\vec{x}(t-1)+\alpha\tanh(\vec{W}_{in}\vec{u}(t)+\bm{\theta}+\vec{\hat{W}}\vec{x}(t-1)),$ (3)
-where $\vec{x}(t)\in R^{n}$ denotes the reservoir state at time $t$, $n$ represents the dimensionality of the reservoir state vector.
-$\vec{W}_{in}\in R^{c\times n}$ is the input-to-reservoir weight matrix, $\bm{\theta}\in R^{n}$ is the bias-to-reservoir weight vector, $\vec{\hat{W}}\in R^{n\times n}$ is the recurrent reservoir weight matrix, $\tanh(.)$ is the element-wise applied hyperbolic tangent activation function, and $\alpha\in[0,1]$ is the leaky parameter, i.e., the decay rate of the nodes.
-The reservoir parameters are initialized according to the constraints specified by the Echo State Property (ESP) and then are left untrained Gallicchio & Micheli (2017).
-Accordingly, the weight values in $\vec{W}_{in}$ and in $\bm{\theta}$ are chosen from a uniform distribution over $[-{scale}_{in},{scale}_{in}]$, where ${scale}_{in}$ represents an input-scaling parameter.
-Matrix $\vec{\hat{W}}$ contains randomly selected values from a uniform distribution.
-It is then adjusted so that the spectral radius of matrix $\vec{\tilde{W}}=(1-\alpha)I+\alpha\vec{\hat{W}}$, denoted as $\rho$, remains below 1 ($\rho<1$).
-This process helps ensure stable dynamics in the reservoir system.
-The reservoir weight $\vec{\hat{W}}$ and input weight $\vec{W_{in}}$ are fixed and random, where each weight is drawn according to Gaussian distribution with parameterized variances.
+The Echo State Network (ESN) (Jaeger and Haas, 2004) implements discrete-time dynamical systems as one of the reservoir computing frameworks (Lukoševičius and Jaeger, 2009; Verstraeten et al., 2007).
+The recurrent reservoir layers are untrained and provide a suffix-based Markovian representation of the past input history (Tiňo et al., 2007), attaching a trained linear readout.
+We use the Leaky Integrator ESN (LI-ESN) model (Jaeger et al., 2007), a variant of the basic ESN, in which leaky integrator reservoir units are adopted.
+In a LI-ESN with $N_{u}$ input units and $N_{r}$ reservoir units the state is updated according to the following state transition function:
+$\mathbf{x}_{t}=(1-\alpha)\mathbf{x}_{t-1}+\alpha\tanh(\mathbf{W}_{in}\mathbf{h}_{t}+\boldsymbol{\theta}+\mathbf{W}\mathbf{x}_{t-1}),$ (4)
+where $\mathbf{x}_{t}\in\mathbb{R}^{N_{r}}$ denote the input and the reservoir state at time $t$, respectively.
+$\mathbf{W}_{in}\in\mathbb{R}^{N_{r}\times d_{\epsilon}}$ is the input-to-reservoir weight matrix, $\boldsymbol{\theta}\in\mathbb{R}^{N_{r}}$ is the bias-to-reservoir weight vector.
+$\mathbf{W}\in\mathbb{R}^{N_{r}\times N_{r}}$ is the recurrent reservoir weight matrix, $\tanh$ is the element-wise hyperbolic tangent activation function, and $\alpha\in[0,1]$ is the leaky parameter.
+The reservoir parameters are initialized according to the constraints specified by the Echo State Property (ESP) (Jaeger and Haas, 2004; Tiňo et al., 2007) and then are left untrained, which makes its computing efficient.
+Accordingly, the weight values in $\mathbf{W}_{in}$ and in $\boldsymbol{\theta}$ are chosen from a uniform distribution over $[-\sigma_{in},\sigma_{in}]$, where $\sigma_{in}$ represents an input-scaling parameter.
+The values in matrix $\mathbf{W}$ are randomly selected from a uniform distribution and then re-scaled such that the spectral radius (i.e.
+the largest eigenvalue in absolute value) of matrix $\mathbf{W}=(1-\alpha)I+\alpha\mathbf{W}$, denoted by $\rho$, is smaller than 1, i.e.
+$\rho<1$ (Jaeger and Haas, 2004).
+Note that the leaky parameter $\alpha$ and the spectral radius of the recurrent weight matrix $\rho$ represent two relevant reservoir computing hyper-parameters.
+$\alpha$ is set to a fixed value of $0.7$ for simplicity, and $\rho$ is uniform randomly sampled between [00,$1$].
 3.3 Linear Readout
 As reservoir output, a readout component is used to linearly combine the outputs of all the reservoir units as in a standard ESN. The output of a reservoir at each time step $t$ can be computed as
-$y(t)=\vec{W}_{out}{\vec{x}(t)}^{T}+\theta_{out},$ (4)
-where $y(t)\in R^{m}$ is the linear readout with $m$ dimensionality.
-$\vec{W}_{out}\in R^{n\times m}$ represents the reservoir-to-readout weight matrix, connecting the reservoir units to the units in the readout.
-Since only $\vec{W}_{out}$ are trained, the optimization problem boils down to linear regression.
-Training is typically not a limiting factor in DRC, in sharp contrast with other neural network architectures.
-The expressiveness and power of reservoir computing rather lie in the high-dimensional non-linear dynamics of the reservoir.
-The reservoir is effective for managing lengthy inputs.
-When we combine it with a Transformer model, the time complexity becomes $\mathcal{O}((k+m)^{2}\times c)$.
-For the Transformer, if we include more time steps in the input, the training time increases quadratically with the input length.
-On the other hand, the reservoir treats each time step individually, keeping the entire history in its learning process without making the input longer.
-This makes the reservoir a better choice for handling long sequences of input data.
-2.3 Baseline Transformer
-In Transformer, the architecture unravels the temporal dependencies into individual time steps (Wu et al., 2020).
-At each time of training, we optimize $\phi$ parameters by a Transformer $\mathcal{M}$ to learn the distribution $Pr(\vec{y}(t+1)|\vec{u}(t-s:t+1),\vec{y}(t-s:t))$ composed of two inputs: the current time stamp $\vec{u}(t+1)$ and that of the previous $s$ time steps $\vec{u}(t-s:t),\vec{y}(t-s:t)$:
-$\vec{g}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)(\vec{u}(t-s:t),\vec{y}(t-s:t))$ (5)
-Here $\kappa$ is a learning parameter adjusting the weights of the current input and previous inputs.
-Thus $\vec{g}(t+1)$ is the input to Transformer, $\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{g}(t+1))$.
-Nonetheless, employing $\vec{g}(t+1)$ as the input for the Transformer poses two significant challenges.
-Firstly, the Transformer’s input length is constrained by a factorization approach involving a retrospective window of size $s$, thereby increasing computational complexity.
-Secondly, the Transformer imposes restrictions on input length, preventing it from exceeding the value of $s$ due to quadratic time complexity considerations.
-Consequently, the incorporation of extensive historical data becomes impractical, limiting its potential impact on the learning and decision-making processes.
-3 Method
-3.1 Readout with Self-Attention
-One downside of the reservoir is that if the reservoir output size $m$ is large, the Transformer training is slow due to its quadratic complexity.
-Since a linear readout of a reservoir needs a much larger size to have the same expressive power as the nonlinear readout, we introduce nonlinear readout layers, which work quite well in reducing DRC output dimension and improving prediction performance.
-Replacing linear readout with nonlinear readout leads to faster convergence and reduces the possibility of getting stuck in a local optimum (Triefenbach & Martens, 2011).
-To model nonlinear readout, we use attention mechanisms (Vaswani et al., 2017) to capture input feature importance, alleviating the problem of vanishing gradient of long-distance dependency.
-There have been various implementations of attention mechanisms.
-We implement our attention model as in (Zhao, 2022).
-Attention mechanisms process sequential data that considers the context for each timestamp.
-$W$ and $b$ are this representation’s weight and bias, and the hyperbolic tangent function $\tanh(.)$ is a nonlinear activation.
-$h_{i,j}=\tanh(\vec{x}(t)_{i}^{T}\vec{W_{t}}+\vec{x}(t)_{j}^{T}\vec{W_{x}}+b_{i})$ (6)
-$e_{i,j}=\sigma(\vec{W_{\alpha}}h_{i,j}+b_{\alpha})$ (7)
-$\gamma_{i,j}=\frac{exp(e_{i,j})}{\sum_{i=1}^{J}(exp(e_{i,j}))}$ (8)
-$r(t)=\left\{\sum_{j}\gamma_{i,j}\vec{x}(t)_{j}\right\}_{i=1}^{c}$ (9)
-Here, $r(t)$ is a non-linear readout using the attention mechanism and the subscript $i$ and $j$ are the $i$-th and $j$-th components of $\vec{x}(t)$.
-The attention weight $\gamma$ is a softmax of $e$.
+$\mathbf{y}_{t}=\mathbf{W}_{out}{\mathbf{x}_{t}}+\boldsymbol{\theta}_{out},$ (5)
+where $\mathbf{y}_{t}\in\mathbb{R}^{m}$ is the linear readout with $m$ dimensionality.
+$\mathbf{W}_{out}\in\mathbb{R}^{m\times N_{r}}$ represents the reservoir-to-readout weight matrix, connecting the reservoir units to the units in the readout.
+$\boldsymbol{\theta}_{out}\in\mathbb{R}^{m}$ is the bias-to-readout weight vector.
+Training the readout involves solving a linear regression problem, typically approached by using Moore-Penrose pseudo-inversion, with further details in  Lukoševičius and Jaeger (2009); Jaeger and Haas (2004).
+Since only $\mathbf{W}_{out}$ are trained, the optimization problem boils down to linear regression.
+3.4 Nonlinear Readout
+The output of the reservoir is the input of the Transformer.
+One downside of the reservoir is that if the reservoir output size is large, then the Transformer training will be slow due to its quadratic complexity to its input size.
+Since a linear readout of a reservoir needs a much larger size to have the same expressive power as the nonlinear readout, we introduce nonlinear readout layers, which work quite well in reducing reservoir output dimension and improving prediction performance hereafter.
+Therefore, for a faster convergence and avoiding a local optimum (Triefenbach and Martens, 2011), we append a nonlinear readout to the linear readout using the following non-linear activation function:
+$\mathbf{\Lambda}_{t}^{l}=\sigma(\mathbf{y}_{t}^{l}),$ (6)
+where $\sigma(\mathbf{y}_{t}^{l})$ is the $l$-th reservoir output, and we use $L$ many reserviors in the following section.
 3.5 Group Reservoir
-In chaos theory, the butterfly effect is the sensitive dependence on initial conditions in which a small change in one state of a deterministic nonlinear system can result in large differences in a later state.
-To enhance reservoir performance, we integrate multiple reservoirs’ nonlinear readout layers, ensuring their independence.
-We consider $Q$ reservoirs with distinct decay rates ($\alpha$ and $\rho$) initialized randomly Gallicchio et al.
+In Chaos theory, the butterfly effect refers to the sensitive dependence on initial conditions, where a small change in one state of a deterministic nonlinear system can result in large differences in a later state.
+As ensemble reduces the variants of the prediction, we integrate multiple reservoirs’ nonlinear readout layers for the stability of our prediction.
+We consider $L$ reservoirs, each with distinct decay rates ($\alpha$ and $\rho$), initialized randomly (Gallicchio et al., 2017).
+This results in a grouped Echo State Network (ESN) representation $\mathbf{o}_{t}$, achieved by element-wise $\oplus$ addition of all reservoirs’ non-linear activation outputs:
+$\mathbf{o}_{t}=\mathbf{\Lambda}_{t}^{1}\oplus\mathbf{\Lambda}_{t}^{2}\oplus\ldots\oplus\mathbf{\Lambda}_{t}^{L}$ (7)
+3.6 Group Reservoir Readout with Self-Attention
+In order to capture the input feature importance, alleviating the problem of vanishing gradient of long-distance dependency, we have used a self-attention mechanism from the aggregate output of group reservoir in Equation 7 by following Shaw et al.
+(2018) :
+$e_{ij}=\frac{(\mathbf{o}_{t}^{i}W^{Q})(\mathbf{o}_{t}^{j}W^{K})^{T}}{\sqrt{d_{\epsilon}}}$ (8)
+$\alpha_{ij}=\frac{\exp{e_{ij}}}{\sum_{k=1}^{n}\exp{e_{ik}}}$ (9)
+$\mathbf{z_{t}}=\sum_{j=1}^{n}\alpha_{ij}(\mathbf{o}_{t}^{j}W^{V})$ (10)
+3.7 Group Reservoir Transformer (RT)
+Our group reservoir efficiently models observed time-series data in dynamical systems and require minimal training data and computing resources (Gauthier et al., 2021) and is ideal for handling long sequential inputs.
+Here we mainly describe adopting a reservoir for Transformer as one of the most well-known DNN architectures.
+However, our methods are adaptable to other DNN model architectures, for whom we also compared results in our experiments.
+Figure 1: Left: Group Reservoir Transformer; Right: Our RT prediction is closer to the ground truth than the Transformer in Lorenz Extractor.
+Figure 1 (Left) shows the architecture of our reservoir-enhanced Transformer.
+We enhance the non-linear readout layers through self-attention mechanisms within ensemble reservoirs.
+This involves updating these layers and performing element-wise addition of the readouts, resulting in an output denoted as $\mathbf{z}_{t}$.
+This output is concatenated with the present input embedding $\mathbf{h}_{t}$, yielding $\mathbf{f}_{t}$, which then undergoes processing via the Transformer encoder $\mathbb{M}(.)$.
+Our approach enables comprehensive modeling of entire historical data in decision-making.
+While the Transformer captures temporal dependencies, limited by a window of size $w$, our method incorporates reservoir pooling to overcome this limitation.
+This permits the consideration of the entire history without a rise in time complexity.
+By incorporating Equation 7 (ensemble of all $L$ reservoir readouts) and then following Equation 10, we have:
+$\mathbf{f}_{t}=(1-\kappa)\mathbf{z}_{t}\uplus\kappa\mathbf{h}_{t}$ (11)
+$\mathbf{\hat{u}}_{t+1:t+\tau}=\mathbb{M}(\mathbf{f}_{t})$ (12)
+[…]
+3.8 Training
+In this work, we use the Huber loss function (Meyer, 2021) as our objective (Equation 13):
+$\mathcal{L}(\mathbf{u}_{t+1:t+\tau},\mathbf{\hat{u}}_{t+1:t+\tau})=\frac{1}{\tau}\sum_{i=t+1}^{t+\tau}\begin{cases}\frac{1}{2}(\mathbf{u}_{i}-\mathbf{\hat{u}}_{i})^{2},&\text{if }|\mathbf{u}_{i}-\mathbf{\hat{u}}_{i}|\leq\delta\\ \delta(|\mathbf{u}_{i}-\mathbf{\hat{u}}_{i}|-\frac{1}{2}\delta),&\text{otherwise}\end{cases}$ (13)
+Here $\delta$ is an adjustable parameter that controls where the function change occurs to keep the function differentiable.
+For a comprehensive understanding of our training algorithm, refer to the pseudo-code provided in the Appendix B, along with a detailed explanation of the notation employed.
+Initially, we sequentially train the model $\mathbb{M}(.)$ across all $T$ time stamps using all the training dataset.
+Subsequently, we fine-tune our model using the validation set to achieve optimal model performance.
+The model is subjected to testing with varying hyper-parameters such as reservoir size, leaky rate, spectral radius, number of reservoirs, learning rate, attention size, Transformer block, and dropout rate.
+The choice of these hyper-parameters may differ depending on the specific dataset with details in Appendix A.
+4 Experiments
+Baselines: We use well-known time-series forecasting models for our baseline comparisons, including Transformer based systems such as DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2022), TimesNet (Wu et al., 2022a), FEDformer (Zhou et al., 2022), Stationary (Liu et al., 2022), ESTformer (Woo et al., 2022) , LightTS (Campos et al., 2023), and Autoformer (Chen et al., 2021), and pre-trained LLMs-based systems such as TimeLLM (Jin et al., 2024) and GPT2TS (Zhou et al., 2024).
+The experimental setup follows  (Zeng et al., 2023) for all baseline models.
+Datasets: We evaluate RT on several datasets commonly used for benchmarking long-term multivariate forecasting models, including four ETT datasets (ETTh1, ETTh2, ETTm1, and ETTm2) (Zhou et al., 2021), Weather, Electricity (ECL), Traffic, and ILI from Zeng et al.
 (2022).
-Utilizing Equations 3 and 9, if $r^{l}(t)$ represents the nonlinear readout from the attention mechanism of the $l$-th reservoir within $Q$, we combine various leaky parameters ($\alpha$) and spectral radius ($\rho$) across the reservoirs.
-This results $\vec{o}(t)$ in a grouped ESN representation in element-wise $\oplus$ addition to all reservoirs’ attention outputs:
-$\vec{o}(t)=r^{1}(t)\oplus r^{2}(t)\oplus...\oplus r^{Q}(t)$ (10)
-3.3 Reservoir Transformer (RT)
-Our group reservoir efficiently models observed time-series data in dynamical systems and require minimal training data and computing resources (Gauthier et al., 2021) and is ideal for handling long sequential inputs.
-Here we mainly describe adopting a reservoir for Transformer Vaswani et al.
-(2017) as one of the most well-known and valuable DNN architectures.
-However, our methods are adaptable to other DNN model architectures, for whom we also compared results in our experiments.
-Figure a: Non-linear readout for RT.
-Figure a: Non-linear readout for RT.
-Figure b: Ensemble of multiple reservoirs
-Figure 1: Description of deep Reservoir Transformer: In Figure (a), the readout obtained from self-attention, combined with the present input $\vec{u}(t+1)$, serves as the input for training the transformer architecture.
-In Figure (b), the input $\vec{u}(t)$ is processed individually by $Q$ reservoirs.
-The nonlinear output of each reservoir is later combined to form the ultimate output, labeled as $\vec{o}(t)$.
-Figure 1 shows the architecture of our reservoir-enhanced Transformer.
-The left figure shows the ensemble of reservoirs, and the right figure shows how to use a non-linear readout reservoir to help Transformer handle arbitrarily long inputs.
-Our approach enables comprehensive modeling of entire historical data in decision-making.
-While the Transformer captures temporal dependencies (Equation 1), limited by a window of size $s$, our method incorporates reservoir pooling to overcome this limitation.
-This permits the consideration of the entire history without a rise in time complexity.
-By incorporating Equation 10 (ensemble of all $Q$ reservoir readouts), we arrive at:
-$\vec{z}(t+1)=\kappa\vec{u}(t+1)+(1-\kappa)\vec{o}(t)$ (11)
-$\bar{\vec{y}}(t+1)=\mathcal{M}(\vec{z}(t+1))$ (12)
-$\vec{z}(t+1)$ denotes the Transformer $\mathcal{M}$ input with $\kappa$ as a learnable parameter indicating the weights of current multivariate states and reservoir state to predict $\bar{\vec{y}}(t+1)$ by Equation 12.
-In the context of multivariate time series prediction, the target label for forecasting up to $q$ future time steps is denoted as $\vec{u}(t+1:t+1+q;v)$, which is contingent upon the observed series $\vec{u}(1:t)$.
-In the multivariate case, all available features are taken into account with the conditional distribution $Pr(\vec{u}(t+1:t+1+q)|\vec{u}(1:t);\phi)$.
-3.4 Training
-We enhance the non-linear readout layers through self-attention mechanisms within ensemble reservoirs.
-This involves updating these layers and performing element-wise addition of the readouts, resulting in an output denoted as $\vec{o}(t)$.
-This output is concatenated with the present input features $\vec{u}(t+1)$, yielding $z(t+1)$, which then undergoes processing via the Transformer encoder $M(.)$.
-Initially, we sequentially train the model $M(.)$ across all time stamps $1:T$ using the training dataset.
-Subsequently, we fine-tune our models using the validation set to achieve optimal model performance.
-The model is subjected to testing with varying hyper-parameters such as reservoir size, leaky rate, spectral radius, number of reservoirs, learning rate, attention size, Transformer block, and dropout rate.
-The choice of these hyper-parameters may differ depending on the specific dataset employed (details in the appendix).
-For a comprehensive understanding of our training algorithm, refer to the pseudo-code provided in the appendix, along with a detailed explanation of the notation employed.
-In this work, we use the Huber loss function Meyer (2021) as our objective (Equation 13):
-$\mathcal{L}(\vec{y},\bar{\vec{y}})=\frac{1}{T}\sum_{i=1}^{T}\begin{cases}\frac{1}{2}(\vec{y}_{i}-\bar{\vec{y}_{i}})^{2},&\text{if}\ |\vec{y}_{i}-\bar{\vec{y}}|\leq\delta\\ \delta|\vec{y}_{i}-\bar{\vec{y}_{i}}|-\frac{1}{2}\delta,&\text{otherwise},\end{cases}$ (13)
-Here $\delta$ is an adjustable parameter that controls where the function change occurs to keep the function differentiable, $T$ is the total number of samples and $\vec{y},\bar{\vec{y}}$ are ground truth and predicted values respectively.
-In the same way, for time series
-4 Experiments
-We show our experimental results of Reservoir Transformer (RT) compared to baselines, including state-of-the-art methods in time-series prediction, such as: Nlinear (Zeng et al., 2022), FEDformer (Zhou et al., 2022), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), LogTrans (Li et al., 2019), GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), STMVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017), and Transformer Vaswani et al.
-(2017).
-To conduct experiments on multivariate time series, we adhered to the established methodology outlined by  (Zeng et al., 2022), which includes data splitting, preprocessing, and setting horizon times.”
-4.1 Result Analysis
-We thoroughly assess our RT method on different time series regression tasks.
-These tasks include Electricity, Traffic, Weather, ETTh1, ETTh2, ETTm1, ETTm2, ILI, Exchange Rate (Zhou et al., 2021), Air Quality (Cini et al., 2021), Daily website visitors (DWV), Daily Gold Price (DGC), and Daily Demand Forecasting Orders (DDFO), Bitcoin Historical Dataset (BTC).
-We also test RT’s performance on two-time series classification tasks, i.e., Absenteeism at work (AW) and Temperature Readings from IOT Devices (details in the Appendix A.1).
-In our empirical evaluation, presented in Table 3, a comparison of multivariate long-term forecasting errors using the Mean Squared Error (MSE) metric was performed across various methods and datasets.
-For each dataset, a forecasting horizon was specified with ILI having horizons $T\in\{24,36,48,60\}$ and the rest having $T\in\{96,192,336,720\}$.
-It is evident that the RT consistently outperforms other methods, achieving the lowest MSE in most scenarios.
-The Transformer-based methods also show competitive performance, with some of their results being underscored for distinction (univariate time forecasting experiments in the Appendix A.5).
-Test Chaotic Behavior:
-Figure a: Mackey Glass Series.
-Figure b: ECG Signal
-Figure a: Mackey Glass Series.
-Figure b: ECG Signal
-Figure c: Lorenz Attractor
-Figure 2: The figures (a), (b), and (c) represent chaotic datasets.
-The blue line illustrates the multi-horizon (800) predictions generated by our model, while the green line depicts the actual ground truth data.
-Time Series $D_{2}$
-Length/Dataset MGS ECG Lorenz MGS ECG Lorenz
-50 0.173 0.301 0.313 3.165 2.824 2.425
-100 0.185 0.314 0.331 3.242 2.913 2.515
-500 0.206 0.341 0.359 3.452 3.173 2.721
-1000 0.225 0.374 0.447 3.531 3.282 2.841
-Table 1: Comparison of Time Series Lengths and Correlation Dimensions ($D_{2}$) for Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets across varying lengths of 50, 100, 500, and 1000.
-Higher values of $D_{2}$ indicate increased complexity in the underlying dynamics.
-In order to investigate the presence of chaotic behavior and assess the predictive capabilities of our RT models, three distinct chaotic datasets were employed: the Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor.
-The experiments involved horizons of 50, 100, 500, and 1000, as outlined in Table 1.
-Our findings indicate that our RT model successfully captures chaotic dynamics across multiple forecasting horizons, shown in Figure 2.
-In addition, to show that the datasets we have used are chaotic, we employ the correlation dimension, denoted as $D_{2}$ serves as a reliable measure of chaotic behavior, with higher values reflecting more pronounced chaos.
-To evaluate the chaotic behavior of various time series datasets  (Pánis et al., 2020), the $D_{2}$ is defined as
-$D_{2}=\underset{\epsilon\to 0}{\lim}\frac{\ln{C(\epsilon)}}{\ln{\epsilon}}.$ (14)
-The correlation sum $C(\epsilon)$ for some small scalar $\epsilon$ is defined as in (Grassberger & Procaccia, 1983):
-$C(\epsilon)=\underset{N\to\infty}{\lim}\frac{2}{N(N-1)}\sum_{i<j}H(\epsilon-|x_{i}-x_{j}|),$ (15)
-where $H$ is the Heaviside step function, $N$ is the number of points, and $|x_{i}-x_{j}|$ is the distance between two points .
-We compute the correlation dimension for all datasets in our experiments using the implementation provided by (Vasquez-Correa, 2015)
-Length/Dataset ETTh1 ETTh2 ETTm1 ETTm2 Exchange ILI BTC DGP DWV DDFO
-100 0.790 1.190 0.560 0.299 0.995 0.995 0.961 1.070 1.343 0.540
-500 1.340 0.973 0.931 0.535 1.099 1.249 0.853 1.626 1.823 0.540
-1000 1.266 0.866 1.061 0.667 1.359 1.299 0.853 1.398 1.838 0.540
-Table 2: This study investigates the correlation dimension $D_{2}$ across multiple datasets using diverse state numbers (100, 1000, and 10000).
-The correlation dimension $D_{2}$ serves as an indicator of chaotic behavior, with higher values denoting a more pronounced presence of chaos (Panis, 2020).
-Our findings indicate that the correlation dimension $D_{2}$ was evaluated for all datasets under varying state numbers.
-Table  1 and  2 display the outcomes of a correlation dimension ($D_{2}$) analysis across various datasets.
-The correlation dimension gauges chaotic tendencies, with higher values suggesting heightened chaos.
-This assessment involves different state numbers: 50, 100, 500, and 1000 and it exhibits $D_{2}$ values for each dataset and state number.
-The findings reveal that, for all datasets, the correlation dimension increases with higher state numbers, indicating intensified chaotic behavior.
-Since Mackey Glass Series (MGS), Electrocardiogram Signal (ECG), and Lorenz Attractor datasets are highly chaotic datasets, they showed the higher $D_{2}$.
-Model
-Horizons
-Linear
-NLinear
+To test the chaotic behavior on these datasets, we employ the correlation dimension, denoted as $D_{2}$ to serve as a reliable measure of chaotic behavior, with higher values reflecting more pronounced chaos (Pánis et al., 2020).
+We compute correlation dimension $D_{2}$ across multiple datasets using diverse state numbers ($100$, $500$, and $1000$).
+Table 1 shows the $D_{2}$ of each dataset.
+Length ETTh1 ETTh2 ETTm1 ETTm2 DDFO Exchange ILI BTC DGP DWV MGS ECG LA
+100 0.790 1.190 0.560 0.299 0.540 0.995 0.995 0.961 1.070 1.343 3.242 2.913 2.515
+500 1.340 0.973 0.931 0.535 0.540 1.099 1.249 0.853 1.626 1.823 3.452 3.173 2.721
+1000 1.266 0.866 1.061 0.667 0.540 1.359 1.299 0.853 1.398 1.838 3.531 3.282 2.841
+low chao medium chao high chao
+Table 1: $D_{2}$ chaotic behavior measure: MGS, ECG, LA are high chaos; ETTh are low chaos.
+Results: The time-series forecasting prediction error rate results are presented in Table 2.
+The input time series length $k$ is set to 300, and we test four different prediction horizons $h\in\{96,192,336,720\}$.
+The evaluation metrics are mean square error (MSE) and mean absolute error (MAE).
+Our RT consistently outperforms all baselines in most cases, achieving significant error reductions.
+Specifically, compared to DLinear, RT reduces MSE by an average of 4.2%.
+Similarly, the reductions in MSE for PatchTST and TimeNet are 1.4% and 2.4%, respectively.
+Additionally, RT outperforms pre-trained LLM-based time series models in Table 3.
+Table 4 shows that RT model outperforms baseline Transformer on various datasets: Air Quality (AQ), Daily Website Visitors (DWV), Daily Gold Price (DGP), Daily Demand Forecasting Orders (DDF), Bitcoin Historical Dataset (BTC), Absenteeism at Work (AW), and Temperature Readings (TR).
+All results are reported in MSE and MAE values, with only AW and TR reporting Accuracy and F-Score metrics.
+Comparisons with more baselines in other evaluation criteria are detailed in Appendix F.
+Methods
+RT
 DLinear
+PatchTST
+TimesNet
 FEDformer
-Autoformer
-Informer
-Pyraformer
-LogTrans
-Repeat
+Stationary
+ETSformer
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.362 0.392 0.376 0.397 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.479
+192 0.396 0.412 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.534 0.504 0.538 0.504
+336 0.427 0.422 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.588 0.535 0.574 0.521
+720 0.441 0.455 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.643 0.616 0.562 0.535
+avg 0.406 0.418 0.422 0.437 0.413 0.430 0.458 0.450 0.440 0.460 0.570 0.537 0.542 0.510
+ETTh2 96 0.262 0.320 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.476 0.458 0.340 0.391
+192 0.331 0.373 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.512 0.493 0.430 0.439
+336 0.358 0.403 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.552 0.551 0.485 0.479
+720 0.369 0.411 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.562 0.560 0.500 0.497
+avg 0.330 0.376 0.431 0.446 0.330 0.379 0.414 0.427 0.439 0.452 0.602 0.543 0.450 0.459
+ETTm1 96 0.272 0.336 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.386 0.398 0.375 0.398
+192 0.310 0.357 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.459 0.444 0.408 0.410
+336 0.351 0.382 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.495 0.464 0.435 0.428
+720 0.380 0.409 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.585 0.516 0.499 0.462
+avg 0.328 0.371 0.357 0.378 0.351 0.380 0.400 0.406 0.448 0.452 0.481 0.456 0.429 0.425
+ETTm2 96 0.163 0.254 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.192 0.274 0.189 0.280
+192 0.218 0.290 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.280 0.339 0.253 0.319
+336 0.271 0.328 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.334 0.361 0.314 0.357
+720 0.356 0.382 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.417 0.413 0.414 0.413
+avg 0.252 0.627 0.267 0.333 0.255 0.315 0.291 0.333 0.305 0.349 0.371 0.306 0.347 0.293
+Weather 96 0.146 0.196 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.173 0.223 0.197 0.281
+192 0.193 0.236 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.245 0.285 0.237 0.312
+336 0.239 0.240 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.321 0.338 0.298 0.353
+720 0.301 0.316 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.288
+avg 0.219 0.247 0.248 0.300 0.225 0.264 0.259 0.287 0.309 0.360 0.288 0.314 0.271 0.334
+Electricity 96 0.130 0.219 0.140 0.237 0.129 0.222 0.168 0.272 0.193 0.308 0.169 0.273 0.187 0.304
+192 0.154 0.243 0.153 0.249 0.157 0.240 0.184 0.289 0.201 0.315 0.182 0.286 0.199 0.315
+336 0.158 0.245 0.169 0.267 0.163 0.259 0.198 0.300 0.214 0.329 0.200 0.304 0.212 0.329
+720 0.190 0.294 0.203 0.301 0.197 0.290 0.220 0.320 0.246 0.355 0.222 0.321 0.233 0.345
+avg 0.158 0.251 0.166 0.263 0.161 0.252 0.192 0.295 0.214 0.327 0.193 0.296 0.208 0.323
+Traffic 96 0.361 0.248 0.410 0.282 0.360 0.249 0.593 0.321 0.587 0.366 0.612 0.338 0.607 0.392
+192 0.376 0.250 0.423 0.287 0.379 0.256 0.617 0.336 0.604 0.373 0.613 0.340 0.621 0.399
+336 0.382 0.268 0.436 0.296 0.392 0.264 0.629 0.336 0.621 0.383 0.618 0.328 0.622 0.396
+720 0.426 0.281 0.466 0.315 0.432 0.286 0.640 0.350 0.626 0.382 0.653 0.355 0.632 0.396
+avg 0.386 0.261 0.433 0.295 0.390 0.263 0.620 0.336 0.610 0.376 0.628 0.379 0.624 0.340
+ILI 24 1.301 0.739 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 2.294 0.945 2.527 1.020
+36 1.409 0.824 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 1.825 0.848 2.615 1.007
+48 1.520 0.807 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.010 0.900 2.359 0.972
+60 1.553 0.852 2.368 1.096 1.520 0.788 2.027 0.928 2.857 1.157 2.178 0.963 2.487 1.016
+avg 1.439 0.805 2.169 1.041 1.443 0.797 2.139 0.931 2.847 1.144 2.077 0.914 2.497 1.004
+Table 2: Multivariate long-term time-series forecasting: lower values indicate better performance. Forecasting horizons are $h\in\{24,36,48,60\}$ for ILI and $h\in\{96,192,336,720\}$ for the rest.
+Methods
 RT
-Dataset Electricity 96 0.140 0.141 0.140 0.193 0.201 0.274 0.386 0.258 1.588 0.140
-192 0.153 0.154 0.153 0.201 0.222 0.296 0.386 0.266 1.595 0.151
-336 0.169 0.171 0.169 0.214 0.231 0.300 0.378 0.280 1.617 0.168
-720 0.203 0.210 0.203 0.246 0.254 0.373 0.376 0.283 1.647 0.201
-Exchange 96 0.082 0.089 0.081 0.148 0.197 0.847 0.376 0.968 0.081 0.101
-192 0.167 0.180 0.157 0.271 0.300 1.204 1.748 1.040 0.167 0.192
-336 0.328 0.331 0.305 0.460 0.509 1.672 1.874 1.659 0.305 0.305
-720 0.964 1.033 0.643 1.195 1.447 2.478 1.943 1.941 0.823 0.586
-Traffic 96 0.410 0.410 0.410 0.587 0.613 0.719 2.085 0.684 2.723 0.391
-192 0.423 0.423 0.423 0.604 0.616 0.696 0.867 0.685 2.756 0.399
-336 0.436 0.435 0.436 0.621 0.622 0.777 0.869 0.734 2.791 0.425
-720 0.466 0.464 0.466 0.626 0.660 0.864 0.881 0.717 2.811 0.442
-Weather 96 0.176 0.182 0.176 0.217 0.266 0.300 0.896 0.458 0.259 0.173
-192 0.218 0.225 0.220 0.276 0.307 0.598 0.622 0.658 0.309 0.217
-336 0.262 0.271 0.265 0.339 0.359 0.578 0.739 0.797 0.377 0.259
-720 0.326 0.338 0.323 0.403 0.419 1.059 1.004 0.869 0.465 0.314
-ILI 24 1.947 1.683 2.215 3.228 3.483 5.764 1.420 4.480 6.587 0.109
-36 2.182 1.703 1.963 2.670 3.103 4.755 7.394 4.799 7.130 0.109
-48 2.256 1.719 2.130 2.622 2.669 4.763 7.551 4.800 6.575 0.111
-60 2.390 1.819 2.368 2.857 2.770 5.264 7.662 5.278 5.893 0.113
-ETTh1 96 0.375 0.374 0.375 0.376 0.449 0.865 0.664 0.878 1.295 0.375
-192 0.418 0.408 0.405 0.420 0.500 1.008 0.790 1.037 1.325 0.411
-336 0.479 0.429 0.439 0.459 0.521 1.107 0.891 1.238 1.323 0.432
-720 0.624 0.440 0.472 0.506 0.514 1.181 0.963 1.135 1.339 0.436
-ETTh2 96 0.288 0.277 0.289 0.346 0.358 3.755 0.645 2.116 0.432 0.228
-192 0.377 0.344 0.383 0.429 0.456 5.602 0.788 4.315 0.534 0.269
-336 0.452 0.357 0.448 0.496 0.482 4.721 0.907 1.124 0.591 0.289
-720 0.698 0.394 0.605 0.463 0.515 3.647 0.963 3.188 0.588 0.306
-ETTm1 96 0.308 0.306 0.299 0.379 0.505 0.672 0.543 0.600 1.214 0.312
-192 0.340 0.349 0.335 0.426 0.553 0.795 0.557 0.837 1.261 0.348
-336 0.376 0.375 0.369 0.445 0.621 1.212 0.754 1.124 1.283 0.372
-720 0.440 0.433 0.425 0.543 0.671 1.166 0.908 1.153 1.319 0.424
-ETTm2 96 0.168 0.167 0.167 0.203 0.255 0.365 0.435 0.768 0.266 0.151
-192 0.232 0.221 0.224 0.269 0.281 0.533 0.730 0.989 0.340 0.187
-336 0.320 0.274 0.281 0.325 0.339 1.363 1.201 1.334 0.412 0.210
-720 0.413 0.368 0.397 0.421 0.433 3.379 3.625 3.048 0.521 0.285
-Table 3: Comparing multivariate long-term forecasting errors using MSE: lower values indicate better performance.
-Specifically, for the ILI dataset, forecasting horizons are $T\in\{24,36,48,60\}$, while for the remaining datasets, they are $T\in\{96,192,336,720\}$.
-Optimal results are emphasized in bold, while the top results from Transformers are underlined for distinction.”
-Horizons 100 500 1500 2500
-Dataset Model MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 NLinear RT 0.362 0.371 0.399 0.403 0.439 0.438 0.479 0.442 0.692 0.602 0.798 0.714 1.683 1.294 2.832 2.104
-ETTh2 NLinear RT 0.283 0.223 0.350 0.310 0.684 0.303 0.591 0.574 0.947 0.561 1.822 0.935 1.361 0.951 2.891 1.821
-ETTm1 NLinear RT 0.302 0.311 0.350 0.352 0.441 0.357 0.433 0.418 0.849 0.739 1.476 1.102 1.153 0.896 2.768 1.562
-ETTm2 NLinear RT 0.164 0.127 0.240 0.217 0.411 0.267 0.432 0.329 1.084 0.740 2.148 0.977 2.099 1.180 3.260 1.936
-Table 4: Comparison of NLinear and RT for explaining the Lyapunov Exponent’s predictive performance across different datasets.
-The table presents MSE and MAE values for both models at various prediction horizons: 100, 500, 1500, and 2500.
-The results demonstrate that the RT model consistently outperforms the NLinear model, achieving significantly lower MSE and MAE values, thus overcoming the Lyapunov Exponent’s challenges of far-sighted forecasting.
-Figure 3: Visualization comparing the prediction of the Lyapunov Exponent using NLinear and RT models. The X-axis corresponds to the time sequence, while the Y-axis corresponds to the target feature.
-4.2 Ablation Study
-Lyapunov Exponent: The Lyapunov Exponent (LE) measures how chaotic a system is by computing how quickly points on a path move away from each other over time.
-Table 4 shows our RT and a baseline NLinear LE over different time steps of 100, 500, 1500, and 2500, in future on different datasets (ETTh1, ETTh2, ETTm1, ETTm2) and outperforms the NLinear model in predicting the LE.
-Figure 3 illustrates the comparison of forecasting 2500 future time steps between the NLinear and RT models, where RT matches more the ground truth than NLinear.
-Entropy Datasets $\vec{u}(t)$ $\vec{o}(t)$ $\vec{z}(t)$ ETTh1 1.011 1.966 2.331 ETTh2 0.993 1.892 2.015 ETTm1 0.816 2.056 2.831 ETTm2 0.599 1.248 2.131 [Uncaptioned image] Table 5: Top: Entropy in different datasets.
-This table presents the entropy values of the original inputs ($\vec{u}(t)$), the corresponding reservoir readouts ($\vec{o}(t)$), and the resultant transformed representations ($\vec{z}(t)$) across various datasets.
-Bottom: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown for ETTh1
-[Uncaptioned image] Figure 4: In the presented graph, the X-axis represents the size of the model parameters in thousands, while the Y-axis illustrates the mean absolute error of the prediction outputs.
-The yellow boxes in the graph indicate the history lengths used in the computation of the RT-1R, RT-5R, and RT-10R models, which were each constructed with one, five, and ten reservoirs, respectively.
-Explain Signature Features: We illustrate and measure how well-distributed the features input of Transformer with or without reservoir are.
-More distribution means higher entropy and more informative, and vice versa.
-We use the Local Interpretable Model-agnostic Explanations (LIME) algorithm (Ribeiro et al., 2016).
-Our goal is to explain the decision-making process within the readout layer by employing LIME to generate a local linear approximation of the model.
-The initial step involves generating perturbed input time series by adding noise to the original $\vec{u(t)}$.
-A simpler interpretable model is then trained using these perturbed inputs along with corresponding output time series $\vec{y}(t)$.
-The complexity of the interpretable model depends on the reservoir model’s intricacy, encompassing both linear and non-linear options.
-Ultimately, we utilize the coefficients of the interpretable model, along with LIME-calculated feature importances, to explain the readout layer’s decision.
-This Table (Top) 5 shows that the readout from the reservoir, $\vec{o}(t)$, has higher entropy than Transformer original input features $\vec{u}(t)$.
-By concatenating both using Equation 11, we achieve the highest entropy.
-Additionally, Table (Bottom) 5 illustrates the Lime explanation for both RT and the baseline Transformer (more ablation in the Appendix A.4).
-History length versus parameter size: Conventional approaches for modeling temporal data require a large number of training parameters, proportional to the window size (s) and the number of features (C).
-This becomes infeasible when considering long historical sequences.
-In contrast, reservoir computing approaches, such as the proposed method, only train a vector representation (i.e.
+TIME-LLM
+GPT2TS
+Methods
+RT
+TIME-LLM
+GPT2TS
+Metric MSE MAE MSE MAE MSE MAE Metric MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.361 0.386 0.362 0.392 0.376 0.397 Weather 96 0.146 0.196 0.147 0.201 0.162 0.212
+192 0.396 0.412 0.398 0.418 0.416 0.418 192 0.193 0.236 0.189 0.234 0.204 0.248
+336 0.427 0.422 0.4427 0.433 0.439 0.433 336 0.239 0.240 0.262 0.279 0.254 0.286
+720 0.441 0.455 0.442 0.457 0.477 0.456 720 0.301 0.316 0.300 0.316 0.326 0.337
+avg 0.408 0.423 0.408 0.423 0.465 0.455 avg 0.219 0.247 0.225 0.257 0.237 0.270
+ETTh2 96 0.262 0.320 0.268 0.328 0.285 0.342 Electricity 96 0.130 0.222 0.131 0.224 0.139 0.238
+192 0.331 0.373 0.329 0.375 0.354 0.389 192 0.154 0.243 0.152 0.241 0.153 0.251
+336 0.358 0.403 0.368 0.409 0.373 0.407 336 0.158 0.245 0.160 0.248 0.169 0.266
+720 0.369 0.411 0.372 0.420 0.406 0.441 720 0.190 0.294 0.192 0.298 0.206 0.297
+avg 0.330 0.376 0.334 0.383 0.381 0.412 avg 0.158 0.251 0.158 0.252 0.167 0.263
+ETTm1 96 0.272 0.336 0.272 0.334 0.292 0.346 Traffic 96 0.361 0.248 0.362 0.248 0.388 0.282
+192 0.310 0.357 0.310 0.358 0.332 0.372 192 0.376 0.250 0.374 0.247 0.407 0.290
+336 0.351 0.382 0.352 0.384 0.366 0.394 336 0.382 0.268 0.385 0.271 0.412 0.294
+720 0.380 0.409 0.383 0.411 0.417 0.421 720 0.426 0.281 0.430 0.288 0.450 0.312
+avg 0.328 0.371 0.329 0.372 0.388 0.403 avg 0.386 0.261 0.388 0.264 0.414 0.294
+ETTm2 96 0.163 0.254 0.161 0.253 0.173 0.262 ILI 24 1.301 0.739 1.285 0.727 2.063 0.881
+192 0.218 0.290 0.219 0.293 0.229 0.301 36 1.409 0.824 1.404 0.814 1.868 0.892
+336 0.271 0.328 0.271 0.329 0.286 0.341 48 1.520 0.807 1.523 0.807 1.790 0.884
+720 0.356 0.382 0.352 0.379 0.378 0.401 60 1.529 0.852 1.531 0.854 1.979 0.957
+avg 0.252 0.316 0.251 0.313 0.284 0.339 avg 1.439 0.805 1.435 0.801 1.925 0.903
+Avg. Training Time 3.150 Minutes 55.41 Minutes 14.76 Minutes Avg. Training Time 36.52 Minutes 510.2 Minutes 146.3 Minutes
+Avg. Occupied GPU 2.033 GB 74.42 GB 16.31 GB Avg. Occupied GPU 2.110 GB 77.271 GB 17.680 GB
+Table 3: Multivariate time-series forecasting (low chaos) comparison with pre-trained LLM model.
+‘Avg.
+Training Time’ is the average training time for the model.
+‘Avg.
+Occupied GPU’ is the average GPU memory usage for the model.
+Method Air Quality DWV DGP DDF BTC AW TR
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Acc. F-Score Acc. F-Score
+Transformer 89.53 11.34 42.95 5.181 168.6K 353.3 1014 22.27 79.40K 232.1 0.928 0.940 0.649 0.934
+RT (Ours) 70.58 9.040 6.181 1.893 28.79K 119.2 10.17 410.6 211.6 211.6 0.899 0.901 0.802 0.975
+Table 4: Time series comparison (medium chaos) in MSE, MAE, Accuracy, and F-score with RT and baseline Transformer. ‘Acc.’ is the Accuracy.
+Signature Feature Extraction: Table 6 shows the entropy of extracted features by baseline ($\mathbf{h}_{t}$), reservoir ($\mathbf{z}_{t}$), and RT ($\mathbf{f}_{t}$).
+Our RT model extracts more informative features than the baseline Transformer model.
+Using the Shannon entropy to measure the information amount, our reservoir output features $\mathbf{z}_{t}$ have higher entropy (more informative) than the embedding output features $\mathbf{h}_{t}$ in the baseline Transformer.
+Combining both $\mathbf{f}_{t}$ gives the highest entropy among them, demonstrating more information extraction with our RT.
+Details of the experiments are enclosed in the Appendix G.2.
+Method RT (Ours) DLinear PatchTST
+Horizons MGS ECG LA MGS ECG LA MGS ECG LA
+50 0.173 0.301 0.313 0.204 0.321 0.356 0.179 0.297 0.310
+100 0.185 0.314 0.331 0.225 0.343 0.370 0.186 0.315 0.313
+500 0.206 0.341 0.359 0.240 0.371 0.389 0.209 0.347 0.360
+1000 0.225 0.374 0.447 0.250 0.390 0.529 0.228 0.378 0.449
+Table 5: MSE on three typical chaos datasets (high chaos).
+Entropy
+Datasets $\mathbf{h}_{t}$ $\mathbf{z}_{t}$ $\mathbf{f}_{t}$
+ETTh1 1.011 1.966 2.331
+ETTh2 0.993 1.892 2.015
+ETTm1 0.816 2.056 2.831
+ETTm2 0.599 1.248 2.131
+Table 6: Feature entropy.
+4.1 Ablation Analysis
+Mackey Glass Series.
+ECG Signal.
+Air Quality.
+Figure 2: The figures represent chaotic dataset examples. The cyan line is ground truth; the red line is the prediction using RT; and the blue line is the prediction using Transformer.
+Output Examples: Figure 1 (Right) and Figure 2 plots the examples from three datasets. We show that our RT method output is more close to the ground truth than the Transformer baseline.
+Figure 3: MSE vs. reservoir numbers.
+Figure 4: RT outperforms baselines with small parameter size and loss in training.
+Impact of Reservoir Count on Performance: Figure 3 illustrates the relationship between the number of ensembling reservoirs and system performance.
+Increasing the number of reservoirs improves the performance, leading to a reduction in loss until the loss converges for $10$ reservoirs.
+Model Parameter Size versus Loss: Conventional models require a large number of training parameters for long historical sequences.
+In contrast, RT only train a vector representation (i.e.
 readout) for the next time step without explicit conditioning on the entire history.
-Empirical results, illustrated in Figure 4, demonstrate that RT provides a better trade-off between the number of training parameters and loss.
-Impact of Reservoir Initial Conditions
-Figure 5: Comparison of MSE loss across five reservoir initialization techniques: random initialization, zero initialization, constant initialization, normal distribution initialization, and uniform distribution initialization.
-The x-axis represents the initialization methods, while the y-axis denotes the corresponding MSE loss values.
-In this study, we delved into the intricate dynamics of reservoir initialization and its consequential impact on the performance of our RT models.
-The reservoir initialization process lays the foundation for the system’s dynamic behavior, influencing the trajectory of internal states and subsequent information processing capabilities.
-We investigated the impact of different initialization conditions on the performance of our RT model.
-We employed five distinct initialization approaches, both in ensemble configurations with group reservoirs and without group reservoirs, as depicted in Figure 5.
-Firstly, we utilized random initialization, which initializes the reservoir randomly within the range of 0 to 1.
-Subsequently, we employed zero initialization, which sets all reservoir values to zero.
-We also explored constant initialization, where a constant value is randomly selected for initialization.
-Furthermore, we investigated initialization using normal and uniform distributions.
-Our experimentation revealed that among these initialization methods, zero initialization consistently yielded superior results in terms of MSE loss, both in ensemble setups with and without group reservoirs.
-Notably, the ensemble technique with group reservoirs demonstrated the lowest loss across both the ETTh2 and ETTm2 datasets.
+Figure 4 illustrates that RT provides a better trade-off between the number of training parameters and loss.
+Initialization ETTh1 ETTh2 ETTm1 ETTm2 Weather Traffic
+NG G NG G NG G NG G NG G NG G
+Random 0.543 0.402 0.435 0.342 0.447 0.328 0.301 0.228 0.334 0.198 0.424 0.380
+Zero 0.429 0.396 0.368 0.331 0.348 0.310 0.236 0.218 0.215 0.193 0.380 0.376
+Constant 0.673 0.405 0.420 0.333 0.516 0.319 0.265 0.225 0.204 0.196 0.424 0.378
+Normal 0.529 0.401 0.445 0.335 0.435 0.332 0.274 0.236 0.242 0.198 0.452 0.379
+Uniform 0.432 0.399 0.395 0.331 0.401 0.314 0.232 0.219 0.200 0.195 0.401 0.375
+[…]
+Table 7: MSE on different initialization using single reservoir (NG) and group reservoirs (G). Group reservoir eliminates the prediction error sensitivity on the initialization.
+Model Memory (Cache) Footprint Context Length Model Time Complexity Seq. Ops.
+Transformer-XL $(d_{key}+d_{value})\cdot H\cdot k\cdot ly$ $k\cdot ly$ Transformer $O(T^{2}\cdot d_{\epsilon})$ $O(1)$
+Compressive Transformer $d_{\epsilon}\cdot(c+T)\cdot ly$ $(c\cdot r+k)\cdot ly$ PatchTST $O((T/P)^{2}\cdot d_{\epsilon})$ $O(1)$
+Memorizing Transformers $(d_{key}+d_{value})\cdot H\cdot k\cdot T$ T Informer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
+RMT $d_{\epsilon}\cdot p\cdot ly\cdot 2$ T Autoformer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
+AutoCompressors $d_{\epsilon}\cdot p\cdot(v+1)\cdot ly$ T Reformer $O(T\cdot log(T)\cdot d_{\epsilon})$ $O(1)$
+Infini-Transformers $d_{key}\cdot(d_{value}+1)\cdot H\cdot ly$ T RNN $O(T\cdot d_{\epsilon}^{2})$ $O(T)$
+Reservoir $N_{r}^{2}$ T Reservoir $O(T\cdot N_{r}^{2})$ $O(T)$
+RT (Ours) $N_{r}^{2}\cdot L$ T RT (Ours) $O(T\cdot k^{2}\cdot d_{\epsilon})$ $O(T)$
+Table 8: Comparison of memory (cache) footprint and time complexity.
+Initialization Comparison: Table 7 shows how different initialization affects the performance for each of the datasets.
+We compute the the $p$-value to see whether the variation in performance due to different initialization is significant when comparing a single reservoir (Non-Group denoted with ‘NG’) with an group reservoir (Group denoted with ‘G’).
+We use the combined mean for each dataset as the population mean to compute the $p$-values.
+We observe significant differences for all the datasets for only the Non-Group experiments.
+Complexity Comparison: Table 8 shows the memory (cache) footprint and time complexity of different models.
+‘Seq.
+Ops.’ is the ‘Sequential Operations’, and the notations are as follows.
+$N_{r}$ is the reservoir state; $L$ is the number of reservoirs in the group reservoir; $k$ is the short-term context window length; $T$ is the long-term context total input length; $ly$ is the number of layers; $H$ is the number of attention heads; $c$ is the Compressive Transformer memory size; $r$ is the compression ratio; $p$ is the number of soft-prompt summary vectors; $v$ is the summary vector accumulation steps; $s$ is the kernel size; $P$ is the patch size; $d_{\epsilon}$ is the embedding model dimension; $d_{key}$ is the dimension of keys in attention; and $d_{value}$ is the dimension of values in attention.
+RT maintains a constant memory complexity of $N_{r}^{2}\cdot L$, and the time complexity is linear regarding total input length.
+Additionally, RT updates its memory with each new input, resulting in retrieval and update occurring once per input.
+Other methods like Transformer-XL (Dai et al., 2019), Compressive Transformer (Rae et al., 2019), Memorizing Transformers (Wu et al., 2022b), RMT (Bulatov et al., 2022), and AutoCompressors (Chevalier et al., 2023) do not have constant memory usage; it depends on the total input segment length $N$ and still exhibits a large context-memory complexity.
+Furthermore, RMT (Bulatov et al., 2022) and AutoCompressors (Chevalier et al., 2023) allow for a potentially infinite context length since they compress the input into summary vectors and then pass them as extra soft-prompt inputs for the subsequent segments.
+However, in practice, the success of those techniques highly depends on the size of soft-prompt vectors (Munkhdalai et al., 2024).
 5 Related Work
-Chaotic time-series prediction using machine learning approaches has been investigated in the last decade using CNN, WNN, FNN, LSTM, etc.
-(Ramadevi & Bingi, 2022).
-For example,(Sangiorgio & Dercole, 2020) proposed the use of a Recurrent Neural Network (RNN) with Long Short-Term Memory (LSTM) cells for chaotic time-series prediction and demonstrated improved performance on several chaotic systems.(Karasu & Altan, 2022) showed a new type of LSTM cell to handle chaotic time series prediction and reported that it outperforms traditional LSTM on several chaotic systems.
-Despite the domain knowledge studied in broad areas, developing new machine learning technologies for Chao’s purpose still needs more attention.
-In particular, the connection between chaotic prediction and conventional time-series forecasting is yet to be more closely integrated (Sahoo et al., 2019).
-There has been extensive literature on time-series forecasting as well  (PaperwithCodes, 2022), and sometimes these datasets hold chaotic properties as well, including using Transformer- and non-Transformer-based solutions for long-term time series forecasting (LTSF).
-For example,  (Mohammadi Farsani & Pazouki, 2020) showed the use of self-attention in the Transformer model for time-series prediction tasks that improved performance.
-Similarly,  (Huang et al., 2019) proposed the use of self-attention in a Recurrent Neural Network (RNN) architecture for time-series prediction, achieving improved performance on several datasets.
-However, these approaches still need more effective models for long-sequential input handling to consider the long history of the past.
-We have proposed a reservoir Transformer (RT) to reach this aim.
-While reservoirs have found extensive application in time-series tasks, their performance often lags behind the current state-of-the-art outcomes (Gauthier, 2021).
-A concept of Reservoir Transformer  (Shen et al., 2020) was introduced by interspersing the reservoir layers into the regular Transformer layers and showing improvements in wall-clock compute time.
-However, such integration will reduce the feature extraction power of both reservoirs in the long-sequence and Transformer in short sequences.
+Time-series prediction using machine learning approaches has been investigated in the last decade using CNN, WNN, FNN, LSTM, etc.
+(Ramadevi and Bingi, 2022; Sangiorgio and Dercole, 2020; Karasu and Altan, 2022; Sahoo et al., 2019; PaperwithCodes, 2022) on datasets with chaotic properties.
+Recently, Transformer-based solutions have shown success for long-term time series forecasting (LTSF) (Mohammadi Farsani and Pazouki, 2020; Huang et al., 2019).
+Recently, researchers have also used various methods to improve long-term time series forecasting  (Zhou et al., 2022, 2021; Wu et al., 2021).
+Nonetheless, the Transformer is expensive in terms of quadratic time and space complexity of the input length.
+Important works have been proposed in improving Transformer architectures and their complexity in long context in various ways, however, compressing the long input sequence into fixed length often misses the dependency of the temporal events.
+Researchers have also shown how using pre-trained LLMs for time-series data improves performance compared to training from scratch (Jin et al., 2023; Chang et al., 2024; Huang, 2024; Munkhdalai et al., 2024).
+de Zarzà et al.
+(2023) used a combination of LLMs (LLaMA-2 and Zephyr-7b-$\alpha$) along with Multimodal LLM (LLaVA) for time-series forecasting.
+However, there is still a restriction on the input length due to the complexity of the neural network models.
+Note that we will experimentally compare with (Munkhdalai et al., 2024) once their code releases.
+Reservoir networks with their simplified training and superior performance for time-series forecasting compared to the traditional models, including time-series kernels, feed-forward networks, and RNN models, has been well documented.
+Researchers have also compared different types of reservoir networks for chaotic time-series data (Bo et al., 2020).
+However, the reservoir network fails to effectively handle long context time series data within fixed-size reservoir (Lukoševičius and Jaeger, 2009; Kim and King, 2020; Mansoor et al., 2021; Bianchi et al., 2020).
+We have introduced reservoir computing to solve the input length problem.
+A concept of reservoir Transformer  (Shen et al., 2020; Gauthier et al., 2021) was introduced by interspersing the reservoir layers into the regular Transformer layers and showing improvements in wall-clock compute time.
+These models improve the training efficiency by using parallelization of time-series methods  (Bollt, 2021).
+However, such integration will reduce the feature extraction power of both reservoirs in the long sequences and Transformer in short sequences.
 We do not combine the reservoir and Transformer linearly but in a non-linear way and receive significant improvements over all previous methods in time-series prediction, including chaotic prediction.
-Additionally, there has been research on using ensembles of models for time-series prediction.(Feng et al., 2019) proposed an ensemble method that combines multiple hybrid models using wavelet transform for chaotic time-series prediction and achieved improved performance on several chaotic systems.(Tang et al., 2020) proposed an ensemble method that combines multiple deep-learning models for time-series prediction and showed improved performance on several datasets.
-In our work, we apply ensemble techniques to our Reservoir Transformer for the first time, contributing to further advancements in time-series prediction.
+Additionally, there has been research on using ensemble of models for time-series prediction.
+Researchers have proposed ensemble methods (Feng et al., 2019) that combine multiple hybrid models using wavelet transform for chaotic time-series prediction and achieved improved performance on several chaotic systems.(Tang et al., 2020; Oliveira and Torgo, 2015; Bertsimas and Boussioux, 2023) proposed an ensemble method that combines multiple deep-learning models for time-series prediction and showed improved performance on several datasets.
+In our work, we propose a novel method to ensemble multiple reservoir networks and combine them with Transformer architecture, contributing to improved performance in time-series prediction.
 6 Conclusion
-We introduce nonlinear readout to reservoir computing that significantly enhances the performance of the state-of-the-art DNNs, including Transformer.
-Our deep Reservoir Transformer has the unique advantage of handling any input length and producing robust prediction outputs.
-Thus, it is ideal for chaotic time-series forecasting.
-This research will open new research questions on seamlessly integrating chaotic analysis, DNNs, and reservoir computing to better understand and predict our environments.
-Appendix A Appendix
-A.1 Dataset and Parameter Configuration
+Our RT has the unique advantage of handling any input length and producing robust prediction outputs.
+Thus, it is ideal for time-series forecasting, including event prediction in chaos.
+We introduce nonlinear readouts to our Reservoir Transformer that outperform state-of-the-art models in terms of accuracy, keeping reasonable complexity in space and time.
+Furthermore, we testify to long-horizon forecasting and show that RT makes the far-horizon prediction less sensitive to initialization.
+This research direction has the potential for succession works by integrating chaotic analysis, deep neural networks, and reservoir computing to understand and predict our environments better.
+Appendix A Dataset and Parameter Configuration
 In this subsection, we outline the approach taken for dataset partitioning and parameter selection within our research framework.
-Dataset Partitioning: The dataset is divided sequentially into three distinct sets: a training set, a validation set, and a test set.
-Specifically, the first $70\%$ of the data is assigned to the training set, while $10\%$ is allocated to the validation set, and the remaining $20\%$ forms the test set.
-This partitioning strategy enables effective model training, validation, and evaluation.
-Model Parameters: The configuration of model parameters plays a crucial role in our research methodology.
-For the Transformer architecture, we employ four Transformer blocks, each comprising 32 multi-head attention mechanisms with a total of four attentions.
+Dataset Partitioning: For the datasets ETTh1, ETTh2, ETTm1, ETTm2, Traffic, Weather, and ILI, we adopted the data preparation settings outlined by [Zeng et al., 2023].
+For other datasets, we divided the data sequentially into three distinct sets: training (70%), validation (10%), and testing (20%).
+This division strategy facilitates efficient training, validation, and testing of models.
+Tools and Hardware: We have used Pytorch and Huggingface libraries to implement our framework. All the experiments were carried out on a single Nvidia RTX-4090 or Tesla H100 GPU.
+Model Parameters: The configuration of model parameters plays a crucial role in experiments.
+For the Transformer architecture, we employ four Transformer blocks, each comprising 12 multi-head attention mechanisms with a total of 12 attention layers.
 In the case of the Feedforward Neural Network (FNN), we utilize hidden layers containing 64 units each.
@@
 Additionally, dropout with a rate of 0.2 is integrated into the attention network.
-Furthermore, the reservoir state size varies between 20 and 50 units across different ensembling reservoirs.
-To introduce diversity in the ensembling process, sparsity levels range from 0.3 to 0.7 for different reservoirs, accompanied by leaky units with values ranging from 0.2 to 0.6.
-Finally, a learning rate of $1\times 10^{-4}$ is adopted to facilitate the training process and achieve optimal convergence.
+Furthermore, the reservoir state size varies between 200 and 300 units across different ensembling reservoirs.
+To introduce diversity in the ensembling process, spectral radius values range from 0.5 to 0.9 for different reservoirs, accompanied by leaky units with values ranging from 0.2 to 0.6.
+Finally, a learning rate of $1\times 10^{-5}$ is adopted to facilitate the training process and achieve optimal convergence.
 By meticulously configuring these parameters, we aim to attain a comprehensive understanding of the model’s performance across various experimental setups.
@@
 We used 10 ensemble reservoirs for this dataset.
-  • Absenteeism at Work:  [^5] https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work The dataset focusing on absenteeism at work is sourced from the UCI repository, specifically the ”Absenteeism at work” dataset.
+  • Absenteeism at Work:  [^5] https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work The dataset focusing on absenteeism at work is sourced from the UCI repository, specifically the "Absenteeism at work" dataset.
 In our analysis, the dataset has been utilized for classification purposes, specifically targeting the classification of individuals as social drinkers.
@@
 The use of ensemble reservoirs enhances the model’s capacity to distinguish between the different temperature reading sources, thereby facilitating accurate classification.
-A.2 Pseudo Algorithm For Deep Reservoir Transformer
+Appendix B Algorithm for RT
 Algorithm 1 Training algorithm of Deep Reservoir Computing
-$\vec{u}(1:T,1:c),\vec{y}(1:T)$
-Optimize $Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t))$ distribution by learning a model $M(.)$
-$\vec{W}_{in},\hat{\vec{W}}\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir
-$\sigma\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir non-linear output
-$\phi\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Transformer encoder $M(.)$
+$\mathbf{u}_{1:T}$, a Transformer model $\mathbb{M}(.)$, a groups reservoirs $\mathcal{R}(.)$, an embedding function $\epsilon(.)$
+Initialization $\mathbb{M}(.)$ and $\mathcal{R}(.)$ with parameters
 while $epoch<epochs$ do
-     for $i\leftarrow 1$ to $T$ do
-         for $k\leftarrow 1$ to $GroupESN$ do
-              $o^{k}(i)\leftarrow R_{k}(\vec{u}(i),\vec{y}(i);\vec{W}_{in},\hat{\vec{W}},\sigma)$triangleright Non-linear readout from $k-th$ reservoir
-              $o(i)\leftarrow o(i)\oplus o^{k}(i)$triangleright Element-wise addition among all reservoirs
+     for $t\leftarrow k$ to $T$ do
+         $\mathbf{h}_{t}=\epsilon(\mathbf{u}_{t-k:t})$triangleright Embedding the time input(3.1)
+         $\mathbf{z}_{t}=\mathcal{R}(\mathbf{h}_{t})$triangleright Readout from Group reservoir (3.2, 3.3, 3.4, 3.5, 3.6 )
+         $\mathbf{f}_{t}=(1-\kappa)\mathbf{z}_{t}\uplus\kappa\mathbf{h}_{t}$triangleright Concatenation (3.7)
+         $\mathbf{\hat{u}}_{t+1:t+\tau}\leftarrow\mathbb{M}(\mathbf{f}_{t})$
      end for
-         $\vec{z}(i+1)\leftarrow\kappa\vec{u}(i+1)+(1-\kappa)o(i+1)$triangleright Concatenation
-         $\bar{\vec{y}}(i+1)\leftarrow\mathcal{M}(\vec{z}(i+1);\phi)$
-     end for
-     $loss\leftarrow\mathcal{L}(\bar{\vec{y}}(i+1),\vec{y}(i+1))$triangleright Calculating loss for every time step $z$
-     $\phi,\sigma\leftarrow update$triangleright Update parameters
+     $loss\leftarrow\mathcal{L}(\mathbf{u}_{t+1:t+\tau},\mathbf{\hat{u}}_{t+1:t+\tau})$triangleright Calculating loss (3.8)
+     update parameters
 end while
-The proposed algorithm 1 is for training a deep reservoir computing model, with the goal of optimizing the distribution of the target variable, $\vec{y}(t+1)$, given the input variables, $\vec{u}(1:t+1)$, and previous target variables, $\vec{y}(1:t)$.
-The algorithm initializes the weights of the reservoir, $\vec{W}_{in}$ and $\hat{W}$, as well as the weights of the non-linear output, $\sigma$, and the Transformer encoder, $\phi$, as random variables sampled from a normal distribution with mean 0 and standard deviation 1.
-The algorithm then enters an outer loop, where the number of iterations is determined by the number of epochs.
-Within this loop, there is an inner loop that iterates through each time step, $i$, from 1 to $T$.
-For each time step, a group of reservoirs, $GroupESN$, are processed in parallel.
-Each reservoir, $R_{k}$, receives the input, $\vec{u}(i)$, and the previous target variable, $\vec{y}(i)$, as input, and calculates a non-linear readout, $o^{k}(i)$, using the weights $\vec{W}_{in},\hat{\vec{W}}$, and $\sigma$.
-The readouts from all reservoirs are then element-wise added together to produce $o(i)$.
-The input of the next time step, $\vec{u}(i+1)$, is concatenated with the readout from the reservoirs, $o(i)$, using a parameter $\kappa$, to produce $\vec{z}(i+1)$.
-Finally, the Transformer encoder, $\mathcal{M}(.)$, is applied to $\vec{z}(i+1)$ using the parameters $\phi$, to produce the target variable for the next time step, $\vec{y}(1+1)$.
-The loss is calculated for every time step, $z$, by comparing the predicted target variable, $\bar{\vec{y}}(1:T)$, with the true target variable, $\vec{y}(1:T)$, using a loss function, $\mathcal{L}(.)$.
-The parameters, $\phi$ and $\sigma$, are then updated to minimize the loss.
-A.3 Variable Descriptions
+In the pseudo algorithm 1, we describe how to train RT models.
+The algorithm initiates by setting parameters for both the Transformer and the reservoir groups.
+During each training epoch, it processes sequences incrementally from a specified start time $k$ up to the total length $T$.
+Each sequence segment transforms an embedded vector, which is then input into the reservoirs to produce intermediate representations.
+These are subsequently merged with the original embedded input, controlled by a weighting parameter $\kappa$, forming a composite feature vector.
+This vector is fed into the Transformer to predict future sequence values.
+The training cycle completes by calculating loss between the predicted and actual values and updating the model parameters accordingly, aiming to minimize prediction errors in following epochs.
+Appendix C Variable Descriptions
+We have described the used variables and math notations in Table 9.
 Variable Description
-$\mathcal{M}(.)$
-A Transformer or DNN model.
-$\vec{u}$
-Samples with $c$ components.
-$c$
-Number of features in dataset
-$\vec{y}$
-Target features for prediction.
-$\bar{\vec{y}}$
-Prediction value
-$T$
-Entire time span or duration
-$t$
-Represents a specific moment or data point in time
-$R$
-Non-linear readout representation of reservoir
-$\vec{o}$
-Aggregate output from all reservoirs within the group
-$\rho$
-Spectral radius of reservoir
-$Q$
-Number of reservoirs in GroupESN
-$m$
-Reservoir output size
-$\vec{W}_{i}n$
-Input-to-reservoir weight matrix
-${\vec{W}_{out}}$
-Reservoir to readout weight matrix
-${\theta_{out}}$
-Reservoir to readout bias matrix
-$\vec{\hat{W}}$
-Recurrent reservoir weight matrix
-$\vec{x}$
-Reservoir state
-$\kappa$
-Parameter to indicate the weights of current state and reservoir state
-$\phi$
-Learnable parameters of $M(.)$
-$s$
-Window size to look back previous sequence
-$R(.)$
-Deep reservoir computing
-$d$
-Feature dimension of Transformer
-$\vec{g}$
-Input for the baseline Transformer
-$\vec{z}$
-The concatenation of reservoir and current multivariate
-$\mathcal{L}$
-Loss function
-$D_{2}$
-Correlation dimension
-$C(\epsilon)$
-Correlation sum
-$H(.)$
-Heaviside step function
+$\mathbb{M}(.)$ Transformer model
+$\mathcal{R}(.)$ Deep reservoir computing
+$\mathbf{u}_{t}$ Input at $t$ time steps
+$\mathbf{x}_{t}$ Reservoir state at $t$ time steps
+$\mathbf{y}_{t}$ Linear output of reservoir
+$\mathbf{h}_{t}$ Embedding of time steps
+$\mathbf{\Lambda}_{t}$ Non-linear readout
+$\mathbf{z}_{t}$ Self attention of group reservoir
+$\mathbf{o}_{t}$ Ensembling of group reservoir
+$\mathbf{f}_{t}$ Input of Transformer
+$T$ Total time input length
+$k$ Look back window
+$H$ Number of attention heads
+$N_{r}$ Reservoir size
+$N_{u}$ Reservoir input size
+$m$ Reservoir output size
+$L$ Number of reservoirs in Group
+$h$ Forecasting horizons
+$d_{\epsilon}$ Model dimension
+$d_{key}$ Dimension of key in attention
+$d_{value}$ Dimension of key in attention
+$l_{y}$ Number of layers in Transformer
+$c$ Compressive Transformer memory size
+$r$ Compression ratio
+$p$ Number of soft-prompt summary vectors
+$s$ Kernel size
+$P$ Patch size
+$\mathbf{\Phi}$ Transformer’s learnable parameters
+$\mathbf{W}_{q}$ Queries weights
+$\mathbf{W}_{v}$ Values weights
+$\mathbf{W}_{k}$ Keys weights
+$\mathbf{W}_{in}$ Input-to-reservoir weight matrix
+$\mathbf{W}$ Reservoir weight matrix
+$\boldsymbol{\theta}$ Bias-to-reservoir weight
+$\mathbf{W}_{out}$ Reservoir-to-readout weight matrix
+$\boldsymbol{\theta}_{out}$ Bias-to-readout weight
 Table 9: Descriptions of all variables used in this paper
-A.4 Ablation Study
-Figure 6: The graph illustrates the relationship between the number of ensemble reservoirs used and the mean square error (MSE) of the prediction outputs.
-The X-axis represents the number of reservoirs included in the ensemble, while the Y-axis depicts the corresponding MSE values.
-The ’star’ in the graph denotes the minimum MSE achieved.
-A.4.1 Impact of Reservoir Count on Performance:
-Figure 6 illustrates the relationship between the number of ensembling reservoirs and system performance.
-The x-axis depicts the count of reservoirs, while the y-axis portrays the Mean Squared Error (MSE) value.
-The graph discernibly demonstrates that increasing the number of reservoirs positively correlates with performance enhancement, leading to a reduction in loss.
-However, beyond a certain threshold of reservoir count (best hyperparameters of the number of reservoirs), the performance improvement saturates, resulting in a plateau in loss reduction.
-This phenomenon suggests an optimal range for reservoir count in relation to performance enhancement.
-A.4.2 Graphical Explanation of Features
-In Figure 7, we present a visual depiction that offers a comprehensive insight into the interpretability of features derived from both the reservoir output $\vec{o}(t)$ and input feature $\vec{u}(t)$.
-Figure 7: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown.
-The test dataset is represented by ETTh1, ETTh2, ETTm1, and ETTm2.
-The results indicate that current input has a greater impact on feature 6, while the reservoir affects multiple features such as 20, 13, 29, and 7.
-A.5 Experiments on Univariate Time sSeries
-Table 7 shows that RT model outperforms GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), ST-MVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017) on the air quality dataset.
+Appendix D Baselines
+We show our experimental results of Reservoir Transformer (RT) compared to baselines, including state-of-the-art methods in time-series prediction, such as FEDformer [Zhou et al., 2022], Autoformer [Wu et al., 2021], Informer [Zhou et al., 2021], Pyraformer [Liu et al., 2021], log(T)rans [Li et al., 2019], GRIN [Cini et al., 2021], BRITS [Cao et al., 2018], STMVL [Yi et al., 2016], M-RNN [Yoon et al., 2018], ImputeTS [Moritz and Bartz-Beielstein, 2017], and Transformer Vaswani et al.
+[2017].
+To conduct experiments on multivariate and uni-variate time series, we evaluate methods outlined by  [Zeng et al., 2022], which includes data splitting, pre-processing, and parameter setting.
+Appendix E Long Uni-Variate Forecasting
+In Table 10 we have compared the long uni-variate time-series forecasting with state of the art time series method and showed improvement of RT in uni-variate forecasting.
+Methods RT Linear NLinear DLinear FEDFormer-f FEDFormer-w Autoformer Informer log(T)rans
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.050 0.172 0.189 0.359 0.053 0.177 0.056 0.180 0.079 0.215 0.080 0.214 0.071 0.206 0.193 0.377 0.283 0.468
+192 0.064 0.201 0.078 0.212 0.069 0.204 0.071 0.204 0.104 0.245 0.105 0.256 0.114 0.262 0.217 0.395 0.234 0.409
+336 0.081 0.225 0.091 0.237 0.081 0.226 0.098 0.244 0.119 0.270 0.120 0.269 0.107 0.258 0.202 0.381 0.386 0.546
+720 0.097 0.292 0.172 0.340 0.080 0.226 0.189 0.359 0.142 0.299 0.127 0.280 0.126 0.283 0.183 0.355 0.475 0.629
+ETTh2 96 0.127 0.275 0.133 0.283 0.129 0.278 0.131 0.279 0.128 0.271 0.156 0.306 0.153 0.306 0.213 0.373 0.217 0.379
+192 0.168 0.324 0.176 0.330 0.169 0.324 0.176 0.329 0.185 0.330 0.238 0.380 0.204 0.351 0.227 0.387 0.281 0.429
+336 0.192 0.351 0.213 0.371 0.194 0.355 0.209 0.367 0.231 0.378 0.271 0.412 0.246 0.389 0.242 0.401 0.293 0.437
+720 0.369 0.411 0.292 0.440 0.225 0.381 0.276 0.426 0.278 0.420 0.288 0.438 0.268 0.409 0.291 0.439 0.218 0.387
+ETTm1 96 0.024 0.119 0.028 0.125 0.026 0.122 0.028 0.123 0.033 0.140 0.036 0.149 0.056 0.183 0.109 0.277 0.049 0.171
+192 0.037 0.149 0.043 0.154 0.039 0.149 0.045 0.156 0.058 0.186 0.069 0.206 0.081 0.216 0.151 0.310 0.157 0.317
+336 0.050 0.168 0.059 0.180 0.052 0.172 0.061 0.182 0.084 0.231 0.071 0.209 0.076 0.218 0.427 0.591 0.289 0.459
+720 0.072 0.199 0.080 0.211 0.073 0.207 0.080 0.210 0.102 0.250 0.105 0.248 0.110 0.267 0.438 0.586 0.430 0.579
+ETTm2 96 0.062 0.182 0.066 0.189 0.063 0.182 0.063 0.183 0.067 0.198 0.063 0.189 0.065 0.189 0.088 0.225 0.075 0.208
+192 0.092 0.224 0.094 0.230 0.090 0.223 0.092 0.227 0.102 0.245 0.110 0.252 0.118 0.256 0.132 0.283 0.129 0.275
+336 0.115 0.258 0.120 0.263 0.117 0.259 0.119 0.261 0.130 0.279 0.147 0.301 0.154 0.305 0.180 0.336 0.154 0.302
+720 0.356 0.382 0.175 0.320 0.170 0.318 0.175 0.115 0.178 0.325 0.219 0.368 0.182 0.335 0.300 0.435 0.160 0.321
+Table 10: Univariate long-term forecasting results
+Appendix F More results
+Table 11 shows that RT model outperforms GRIN [Cini et al., 2021], BRITS [Cao et al., 2018], ST-MVL [Yi et al., 2016], M-RNN [Yoon et al., 2018], ImputeTS [Moritz and Bartz-Beielstein, 2017] on the air quality dataset.
 On regression and classification tasks, RT consistently outperforms the baseline Transformer with respect to all criteria, achieving significantly lower MSE and MAE values across various datasets like DWV, DGP, DDFO, and BTC.
@@
 Through this analysis, we provide evidence that the RT architecture excels not only in regression tasks but also in classification tasks, outperforming traditional Transformer architectures.
+Additionally, we have shown the comparison of multi-variate time series with other methods such as LightTS [Campos et al., 2023] and Autformer [Wu et al., 2021] in Table 11.
 Air Quality
@@
 Daily website visitors (DWV)
-Baseline 98.451 7.583 0.996 42.954 5.181 0.997 $\sim$13.2k
+Baseline 98.45 7.583 0.996 42.95 5.181 0.997 $\sim$13.2k
 RT 10.400 (-89.43%) 2.114 0.999 6.181 1.893 0.999 $\sim$3.2k
@@
 RT 0.802 0.843 0.728 0.794 0.867 0.820 $\sim$2.3k Test
-Table 7: Assessing the performance of regression on both training and test datasets, we utilize a window size of $s=5$ to observe historical data in baseline Transformers.
+Table 11: Assessing the performance of regression on both training and test datasets, we utilize a window size of $100$ to observe historical data in baseline Transformers and RT.
+Methods
+RT
+DLinear
+PatchTST
+TimesNet
+FEDformer
+Stationary
+ETSformer
+LightTS
+Autoformer
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.362 0.392 0.376 0.397 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.479 0.424 0.432 0.449 0.459
+192 0.396 0.412 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.534 0.504 0.538 0.504 0.475 0.462 0.500 0.482
+336 0.427 0.422 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.588 0.535 0.574 0.521 0.518 0.488 0.521 0.496
+720 0.441 0.455 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.643 0.616 0.562 0.535 0.547 0.533 0.514 0.512
+avg 0.406 0.418 0.422 0.437 0.413 0.430 0.458 0.450 0.440 0.460 0.570 0.537 0.542 0.510 0.491 0.479 0.496 0.487
+ETTh2 96 0.262 0.320 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.476 0.458 0.340 0.391 0.397 0.437 0.346 0.388
+192 0.331 0.373 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.512 0.493 0.430 0.439 0.520 0.504 0.456 0.452
+336 0.358 0.403 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.552 0.551 0.485 0.479 0.626 0.559 0.482 0.486
+720 0.369 0.411 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.562 0.560 0.500 0.497 0.863 0.672 0.515 0.511
+avg 0.330 0.376 0.431 0.446 0.330 0.379 0.414 0.427 0.437 0.449 0.526 0.516 0.439 0.452 0.602 0.543 0.450 0.459
+ETTm1 96 0.272 0.336 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.386 0.398 0.375 0.398 0.374 0.400 0.505 0.475
+192 0.310 0.357 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.459 0.444 0.408 0.410 0.400 0.407 0.553 0.496
+336 0.351 0.382 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.495 0.464 0.435 0.428 0.438 0.438 0.621 0.537
+720 0.380 0.409 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.585 0.516 0.499 0.462 0.527 0.502 0.671 0.561
+avg 0.328 0.371 0.357 0.378 0.351 0.380 0.400 0.406 0.448 0.452 0.481 0.456 0.429 0.425 0.435 0.437 0.588 0.517
+ETTm2 96 0.163 0.254 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.192 0.274 0.189 0.280 0.209 0.308 0.255 0.339
+192 0.218 0.290 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.280 0.339 0.253 0.319 0.311 0.382 0.281 0.340
+336 0.271 0.328 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.334 0.361 0.314 0.357 0.442 0.466 0.339 0.372
+720 0.356 0.382 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.417 0.413 0.414 0.413 0.675 0.587 0.433 0.432
+avg 0.252 0.316 0.267 0.333 0.255 0.315 0.291 0.333 0.305 0.349 0.371 0.306 0.347 0.293 0.342 0.409 0.436 0.327
+Weather 96 0.146 0.196 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.173 0.223 0.197 0.281 0.182 0.242 0.266 0.336
+192 0.193 0.236 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.245 0.285 0.237 0.312 0.227 0.287 0.307 0.367
+336 0.239 0.240 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.321 0.338 0.298 0.353 0.282 0.334 0.359 0.395
+720 0.301 0.316 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.288 0.352 0.386 0.419 0.428
+avg 0.219 0.247 0.248 0.300 0.225 0.264 0.259 0.287 0.309 0.360 0.288 0.314 0.271 0.334 0.261 0.312 0.338 0.382
+Electricity 96 0.130 0.219 0.140 0.237 0.129 0.222 0.168 0.272 0.193 0.308 0.169 0.273 0.187 0.304 0.207 0.307 0.201 0.317
+192 0.154 0.243 0.153 0.249 0.157 0.240 0.184 0.289 0.201 0.315 0.182 0.286 0.199 0.315 0.213 0.316 0.222 0.334
+336 0.158 0.245 0.169 0.267 0.163 0.259 0.198 0.300 0.214 0.329 0.200 0.304 0.212 0.329 0.230 0.333 0.231 0.338
+720 0.190 0.294 0.203 0.301 0.197 0.290 0.220 0.320 0.246 0.355 0.222 0.321 0.233 0.345 0.265 0.360 0.254 0.361
+avg 0.158 0.251 0.166 0.263 0.161 0.252 0.192 0.295 0.214 0.327 0.193 0.296 0.208 0.323 0.229 0.329 0.227 0.338
+Traffic 96 0.361 0.248 0.410 0.282 0.360 0.249 0.593 0.321 0.587 0.366 0.612 0.338 0.607 0.392 0.615 0.391 0.613 0.388
+192 0.376 0.250 0.423 0.287 0.379 0.256 0.617 0.336 0.604 0.373 0.613 0.340 0.621 0.399 0.601 0.382 0.616 0.382
+336 0.382 0.268 0.436 0.296 0.392 0.264 0.629 0.336 0.621 0.383 0.618 0.328 0.622 0.396 0.613 0.386 0.622 0.337
+720 0.426 0.382 0.466 0.315 0.432 0.286 0.640 0.350 0.626 0.382 0.653 0.355 0.632 0.396 0.658 0.407 0.660 0.408
+avg 0.386 0.261 0.433 0.295 0.390 0.263 0.620 0.336 0.610 0.376 0.628 0.379 0.624 0.340 0.621 0.396 0.622 0.392
+ILI 24 1.301 0.739 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 2.294 0.945 2.527 1.020 8.313 2.144 3.483 1.287
+36 1.409 0.824 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 1.825 0.848 2.615 1.007 6.631 1.902 3.103 1.148
+48 1.520 0.807 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.010 0.900 2.359 0.972 7.299 1.982 2.669 1.085
+60 1.553 0.852 2.368 1.096 1.520 0.788 2.027 0.928 2.857 1.157 2.178 0.963 2.487 1.016 7.283 1.985 2.770 1.125
+avg 1.439 0.805 2.169 1.041 1.443 0.797 2.139 0.931 2.847 1.144 2.077 0.914 2.497 1.004 7.382 2.003 3.006 1.161
+Table 12: Multivariate long-term time-series forecasting: lower values indicate better performance. Forecasting horizons are $h\in\{24,36,48,60\}$ for ILI and $h\in\{96,192,336,720\}$ for the rest.
+Appendix G More Ablation Analysis
+G.1 Look-Back Window Size vs Loss
+In Figure 5, we have proved that RT works for the short look-back window better than the Dlinear method, since RT can capture the long-term dependencies using group reservoirs.
+Figure 5: Look back window size vs loss
+G.2 Graphical Explanation of Features
+In Figure 6, we present a visual depiction that offers a comprehensive insight into the interpretability of features derived from both the reservoir output $\mathbf{z}_{t}$ and input feature $\mathbf{u}_{t}$.
+Figure 6: The Lime Explanation effect of current input (A) and reservoir (B) on output prediction is shown.
+The test dataset is represented by ETTh1, ETTh2, ETTm1, and ETTm2.
+The results indicate that current input has a greater impact on feature 6, while the reservoir affects multiple features such as 20, 13, 29, and 7.
+G.3 Non-Linear output
+In Table 13, we present a comparison of the non-linear readout of the reservoir using different activation functions.
+Activation ETTh1 ETTh2 Traffic
+Linear ReLu Tanh Self-Attention Linear ReLu Tanh Self-Attention Linear ReLu Tanh Self-Attention
+MSE 0.538 0.462 0.474 0.441 0.447 0.608 0.495 0.485 0.498 0.428 0.463 0.426
+MAE 0.562 0.470 0.475 0.455 0.348 0.552 0.468 0.478 0.327 0.293 0.314 0.281
+Table 13: Comparison of MSE for different readout activation.
+Appendix H Limitations
+Our method assumes that the compressed memory represents the long context within the time-series data.
+However, in practice memory compression can result in potential information loss.
+Therefore, giving the whole context as input instead of compressing it to a fixed size may result in better performance.
+Additionally, we evaluated our method on standard time-series benchmarks and thus set the maximum horizon value as 720.
+There may exist datasets that have extremely long-term dependencies.
+To capture this dependency in such datasets, we would either have to increase the size of the reservoir’s input.